### v1

# From Estimands to Robust Inference of Treatment Effects in Platform Trials

Abstract
A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner and can accelerate the evaluation of new treatments. However, the flexibility that marks the potential of platform trials also creates inferential challenges. Two key challenges are the precise definition of treatment effects and the robust and efficient inference on these effects. To address these challenges, we first define a clinically meaningful estimand that characterizes the treatment effect as a function of the expected outcomes under two given treatments among concurrently eligible patients. Then, we develop weighting and post-stratification methods for estimation of treatment effects with minimal assumptions. To fully leverage the efficiency potential of data from concurrently eligible patients, we also consider a model-assisted approach for baseline covariate adjustment to gain efficiency while maintaining robustness against model misspecification. We derive and compare asymptotic distributions of proposed estimators in theory and propose robust variance estimators. The proposed estimators are empirically evaluated in a simulation study and illustrated using the SIMPLIFY trial. Our methods are implemented in the R package RobinCID.
Keywords: Covariate adjustment; Eligible and concurrent controls; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
1 Introduction
The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Saville and Berry,, 2016; Woodcock and LaVange,, 2017). As a result, the landscape of clinical trials has been evolving, to make clinical trials faster and more efficient while continuing to provide reliable information on safety and effectiveness (The White House Office of Science and Technology,, 2023).
The platform trial has emerged as an innovative trial design to expedite the evaluation of potentially effective treatments (Berry et al.,, 2015; Saville and Berry,, 2016; Woodcock and LaVange,, 2017; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023). It is a dynamic clinical trial that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner, where new treatments can be added to the platform as they become available, and existing ones can leave the platform based on a finding of futility or efficacy. Such a trial structure not only streamlines operations but also has the potential to improve statistical efficiency by facilitating the sharing of data between different treatment evaluations (Mehta et al.,, 2023). The effectiveness of platform trials in addressing urgent medical challenges quickly is evidenced by the hallmark RECOVERY platform trial, which was designed to rapidly evaluate COVID-19 therapies during the pandemic (Normand,, 2021; Buenconsejo et al.,, 2023). Platform trials are also attractive in the exploratory phase to quickly screen multiple treatments for diseases (Wang and Yee,, 2019) and in rare diseases to enhance the chance of identifying effective treatments (Dhaenens et al.,, 2024). Two examples of platform trials are presented in Section 2 to provide detailed illustrations.
Although platform trials offer several advantages over traditional trials, they also create challenges in statistical inference on treatment effects, as highlighted in a recent Food and Drug Administration draft guidance (FDA,, 2023). One primary challenge is how to define a meaningful treatment effect in a platform trial setting, as multiple treatments may target different biomarker-defined subgroups and enter the trial at different times — a complexity that does not exist in traditional trials. In platform trials, when comparing a control and a treatment of interest, many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, (2023) have recommended using eligible and concurrent controls, which refer to the patients who are eligible to receive the treatment and are concurrently randomized to the control arm during the same enrollment period as the other patients being assigned to the treatment arm. The use of eligible and concurrent controls, or concurrently eligible patients, preserves the integrity of randomized comparisons and inference on treatment effects. In contrast, patients who are either ineligible or nonconcurrent may not be comparable to the treated group due to measured and unmeasured differences (FDA,, 2023). In Section 3, we explicitly present, for the first time in the literature, how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients, which aligns with the ICH E9 , 2019 (R1) estimand framework.
The next challenge is on robust estimation of the defined treatment effects for concurrently eligible patients. Given the adaptive nature of platform trials, where treatments can be added and removed, the available treatment options vary over time. This dynamic, coupled with different eligibility criteria for each treatment, means that the probabilities of treatment assignments can vary over time and patient characteristics — another complexity absent in traditional trials. Thus, naively comparing outcome means between two treatment arms may yield biased results, an example of Simpson’s paradox (Collignon et al.,, 2020; Park et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021).
The literature on statistical methods for platform trials is expanding, though it primarily focuses on model-based techniques for incorporating nonconcurrent controls, which can be sensitive to model misspecification (Saville et al.,, 2022; Bofill Roig et al.,, 2022; Elm et al.,, 2012; Wang et al.,, 2023). Some research has been conducted on statistical methods for concurrently eligible patients. Ventz et al., (2018) considered several adaptive randomization procedures and tests for the hypothesis of no treatment effect. Marschner and Schou, (2022) proposed stratification using enrollment windows, without taking into account other factors related to treatment and randomization. Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows. However, there remains a lack of rigorous investigation on how to make robust inference on treatment effects in platform trials under the same minimal statistical assumptions as those in traditional clinical trials.
To fill this gap, in Section 4, we develop methods for estimating treatment effects for concurrently eligible patients using inverse probability weighting or post-stratification to account for varying treatment assignment probabilities. Additionally, we consider model-assisted approaches for covariate adjustment to gain efficiency while maintaining robustness against model misspecification. In Section 5, we theoretically establish the asymptotic distributions of proposed estimators and compare their asymptotic relative efficiency. We provide robust variance estimators in Section 6. Simulation results and an application of our methods to the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) are in Section 7. Our methods are implemented in the R package RobinCID.
2 Examples of Platform Trials
2.1 A stylistic example
First, we present a stylistic example of a platform trial in Figure 1 to provide a concrete illustration of its key features. This trial comprises five sub-studies, each containing a randomized controlled comparison of several treatments. Initiating or stopping a sub-study is governed by a master protocol. Additionally, each sub-study has an intervention-specific appendix outlining protocol elements specific to that sub-study. Patients are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis. Real trials resembling this structure include Lung-MAP for lung cancer (Herbst et al.,, 2015; Redman et al.,, 2020) and an ongoing trial for chronic pain (ClinicalTrials.gov, NCT05986292).
Figure 1: A stylistic example of a platform trial. The x-axis represents calendar time discretized into windows with different potential enrollment in components of the platform trial. The figure illustrates both different treatment comparisons and different patient subgroups evolving over time.
In Figure 1, each box represents a sub-study along with its status, and each striped block within a box represents the enrollment period for a treatment evaluated in that sub-study. The horizontal axis denotes calendar time. Enrollment windows are defined by changes in the available treatments for enrollment. For instance, the initiation of enrollment for sub-study 3 marks the start of enrollment window 2, within which treatment 4 is compared to treatment 1 as the control among patients with disease subtype A. Likewise, the conclusion of the enrollment period for sub-study 2 defines the start of enrollment period 3. Subsequently, the start of treatment 3 as the new control in place of treatment 1 among subtype A defines the start of enrollment window 4.
The following is a narrative of Figure 1 according to the order of enrollment windows.
Enrollment Window 1. In this example, there are two initial sub-studies. Patients with any disease subtype is eligible to sub-study 1, while sub-study 2 is restricted to patients with subtype A. During enrollment window 1, patients with subtype A are randomly assigned to sub-study 1 or 2, and then randomized to treatment 1 or 2 if they are in sub-study 1 and to treatment 1 or 3 if they are in sub-study 2. Patients with disease subtypes other than A are enrolled in sub-study 1 and then randomized to treatment 1 or 2.
Enrollment Window 2. Sub-study 3 is added to the platform trial when a new intervention treatment 4 becomes available for disease subtype A. Following this addition, a patient with subtype A is randomly assigned to either sub-study 1, 2, or 3, and subsequently randomized to the treatments being assessed within the respective sub-study.
Enrollment Window 3. After sub-study 2 reaches full enrollment, it closes its enrollment. Later during this period, treatment 3 meets the pre-specified success criteria.
Enrollment Window 4. Since treatment 3 has been demonstrated to be superior to treatment 1 for subtype A, it replaces treatment 1 as the control for subtype A. Consequently, treatment 3 replaces treatment 1 in sub-study 3. Moreover, sub-study 1 excludes patients with subtype A, and sub-study 4 is introduced to continue evaluating treatment 2 in comparison to the new control treatment 3 for patients with subtype A.
Enrollment Window 5. Sub-study 5 is added upon the introduction of a new treatment 5 for patients positive for biomarker B. In sub-study 5, treatment 5 is evaluated against treatment 1 for patients without subtype A, or to treatment 3 for patients with subtype A.
Enrollment Window 6. Sub-study 4 has finished its planned enrollment and is no longer accepting new patients.
Enrollment Window 7. A new treatment, treatment 6, has been introduced in sub-study 5 for patients positive for biomarker B. During this enrollment window, treatment 2 in sub-study 4 fails to meet the pre-specified success criteria.
As the platform trial continues, additional sub-studies or treatments may be added as new interventions emerge for one or more targeted groups of patients, while existing sub-studies may be concluded due to either meeting success criteria or demonstrating futility.
As exemplified in Figure 1, platform trials can be very flexible: adding new sub-studies and/or treatments with different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial. Moreover, compared to a traditional clinical trial, it enables potential pooling of data across sub-studies. Specifically, in traditional clinical trials, patients self-select into studies and are randomized into treatment arms, with treatment comparisons conducted solely within each study. In contrast, platform trials enroll patients into a master protocol, randomizing them into eligible sub-studies and subsequently treatment arms, and consequently, patients can be pooled across sub-studies. For example, when comparing treatments 1 and 2, in a platform trial like Figure 1, patients from not only sub-study 1 but also other sub-studies receiving treatment 1 can be included in the analysis, whereas we analyze data within sub-study 1 in a traditional trial. See the simulation results and discussions in Section 7.1 for the efficiency gain.
2.2 A real example: the SIMPLIFY trial
Our second example is a real platform design aliased SIMPLIFY for people with cystic fibrosis (Mayer-Hamblett et al.,, 2023). It contains two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the Cystic Fibrosis Therapeutics Development Network. The primary objective for each study is to determine whether discontinuing either hypertonic saline (HS) or dornase alfa (DA), respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator, elexacaftor plus tezacaftor plus ivacaftor (ETI). The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1). Figure 2 depicts the SIMPLIFY study design.
Figure 2: SIMPLIFY study design. All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups: those receiving HS only, those receiving DA only, and those receiving both therapies. Individuals on a single therapy enter the corresponding study, while those on both are randomized to the HS or DA study, first with a 1:1 probability and then with a 3:1 probability during subsequent enrollment. All individuals in HS study (or DA study) are randomized 1:1 to continue or discontinue HS (or DA).
3 Setup and Estimand
3.1 Randomization procedures in platform trials
Consider a platform trial with $J$ treatments and $n$ patients. For each patient $i$, let $\bm{W}_{i}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A_{i}$ be the treatment assignment indicator that equals $j$ if the patient is assigned to arm $j$, $Y_{i}^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y_{i}=Y_{i}^{(A_{i})}$ be the observed outcome. Throughout, $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, are independent and identically distributed realizations of $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments. The following assumption describes the randomization procedure in platform trials.
Assumption 1 (Randomization).
There exists an observed baseline discrete variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
Assumption 1 highlights a fundamental difference between the randomization procedure used in platform trials versus traditional clinical trials. In traditional trials, the randomization probability is typically a fixed positive constant. However, in platform trials, this probability, $\pi_{j}(Z)$, varies depending on a baseline discrete variable $Z$ and may be zero for some treatments. Notably, $Z$ usually includes the enrollment window by calendar time and some key patient characteristics, as the enrollment window determines which treatments are available for enrollment, and patient characteristics determine the treatments for which a patient is eligible. Hence, $\pi_{j}(Z)$ may be zero if treatment $j$ has not entered the platform trial or has closed its enrollment, or if the patient does not meet the eligibility requirements for treatment $j$. Additionally, non-zero values of $\pi_{j}(Z)$ can differ over time and patient characteristics. Values of $\pi_{j}({Z})$, $j=1,...,J$, are known by design, but $\pi_{j}({Z})$ should be viewed as random due to the randomness of ${Z}$.
The example in Figure 1 illustrates these dynamics. Here, $Z=(Z_{\rm win},Z_{\rm sub},Z_{\rm B+})$ represent the enrollment window by calendar time, disease subtype (A or not A), and biomarker B (positive or negative), respectively, where $Z_{\rm B+}$ does not affect treatment assignment until enrollment window 5. For simplicity, consider randomization with equal probabilities. If patient $i$ with subtype A enters the trial during enrollment window 1 in which only sub-studies 1 and 2 are open, this patient is randomized to either sub-study with equal probability. If randomized to sub-study 1, the patient is further randomized to treatment 1 or 2 with equal probability. If randomized to sub-study 2, the patient is further randomized to treatment 1 or 3 with equal probability. Consequently, for this patient $i$, $\pi_{1}({Z}_{i})=0.5\times 0.5+0.5\times 0.5=0.5$, $\pi_{2}({Z}_{i})=0.5\times 0.5=0.25$, $\pi_{3}({Z}_{i})=0.5\times 0.5=0.25$, with all other $\pi_{j}({Z}_{i})$ values equal to 0. If another patient $i^{\prime}$ without subtype A enters the trial during the same enrollment window 1, then this patient is assigned to sub-study 1 with probability 1. Consequently, $\pi_{1}({Z}_{i^{\prime}})=\pi_{2}({Z}_{i^{\prime}})=0.5$ and all other $\pi_{j}({Z}_{i^{\prime}})$ values are 0. Similarly, if another patient $i{{}^{\prime\prime}}$ with subtype A enters the trial during enrollment window 2 in which sub-studies 1,2,3 are open, this patient has $\pi_{1}(Z_{i^{\prime\prime}})=0.5$, $\pi_{2}(Z_{i^{\prime\prime}})=\pi_{3}(Z_{i^{\prime\prime}})=\pi_{4}(Z_{i^{\prime\prime}})=1/6$, with all other $\pi_{j}({Z}_{i^{\prime\prime}})$ values equal to 0. These cases exemplify how the positivity and values of $\pi_{j}(Z)$ can vary with $Z$.
3.2 Treatment effects for concurrently eligible patients
We first define a clinically meaningful estimand (ICH E9 , 2019, R1) for comparing a given pair of treatments, $j$ and $k$, one of which is typically a control. As discussed in the introduction, many researchers and the FDA have recommended that the comparison be based on concurrently eligible patients, i.e., patients who meet the eligibility criteria for both treatments and are enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment $j$ or $k$. Formally, the concurrently eligible patients for treatments $j$ and $k$ are those with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
We provide two examples to illustrate this concept. The first example is based on Figure 1. When an investigator compares treatments 1 and 3 at the end of sub-study 2, the concurrently eligible patients are all subtype A patients from enrollment windows 1 and 2, which includes all patients in sub-study 2, as well as patients with subtype A in sub-study 1 and all patients in sub-study 3 from the first two windows. Patients without subtype A or enrolled in later windows are not included, as they are either ineligible or nonconcurrent. The second example is based on the SIMPLIFY design in Figure 2. The original analysis by Mayer-Hamblett et al., (2023) calculated the effects of discontinuing vs. continuing HS or DA within each study separately. However, this approach did not utilize all concurrently eligible patients with additional control data that could have been used from participants assigned to continue therapy across either study. Thus, when comparing discontinuing HS versus continuing therapy, for instance, the concurrently eligible patients are all individuals on HS regardless of their DA status, which includes all participants in the HS study, as well as those on both HS and DA in the DA study.
We propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ among concurrently eligible patients:
${\bm{\vartheta}}_{jk}=\left(\begin{array}[]{c}\theta_{jk}\\ \theta_{kj}\end{array}\right)=\left(\begin{array}[]{c}E(Y^{(j)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\\ E(Y^{(k)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\end{array}\right),$ (1)
where both subscripts $j$ and $k$ are used in $\bm{\vartheta}_{jk}$, $\theta_{jk}$, and $\theta_{kj}$ because both $\pi_{j}({Z})$ and $\pi_{k}({Z})$ are involved. In essence, a contrast of the estimand in (1) is an unconditional treatment effect for all concurrently eligible patients, averaging across all eligible patients and over time during which the enrollment window is open. Time can represent many background factors that may be changing although this is not explicitly recognized in estimand specification for traditional clinical trials.
Our defined population of concurrently eligible patients is the largest population that can preserve the integrity of randomization when inferring treatment effects. Individuals outside this group have zero probability of being assigned to treatments $j$ or $k$, or both, due to their ineligibility or the unavailability of treatments, thereby necessitating extrapolation. This is the reason why we do not consider the population means $E(Y^{(j)})$ and $E(Y^{(k)})$ that are defined for the entire platform trial population. Conversely, our estimand in (1) represents the treatment effect for a broader population compared to just those enrolled in a specific sub-study. Our estimand in (1) truly respects the eligibility criteria for the given pair of treatments for comparison, rather than on the subset assigned to a particular sub-study.
Throughout the paper, we use the notation convention that $a_{kj}$ is defined by swapping subscripts $j$ and $k$ in $a_{jk}$, e.g., $\theta_{jk}$ and $\theta_{kj}$ defined in (1). Based on the estimand ${\bm{\vartheta}}_{jk}$ in (1), the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$. Our approach to defining treatment effects is adaptable to situations where population characteristics other than means are considered, or when the treatment effect for a targeted subgroup is of primary interest.
4 Estimation Methods
4.1 Naive method
After we define the estimand ${\bm{\vartheta}}_{jk}$ in (1), the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible patients defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z}),\pi_{k}({Z})$ may still vary over time and patient characteristics, leading to confounding.
A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from patients with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$. While $\bar{Y}_{jk}$ is valid in traditional trials, it is typically biased under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome. Specifically, as shown in the Supplementary Material, as $n\to\infty$,
$\bar{Y}_{jk}\to\frac{E\{\pi_{j}({Z})E(Y^{(j)}\mid{Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}{E\{\pi_{j}({Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}\quad\mbox{in probability}.$ (2)
The limit of $\bar{Y}_{jk}$ in (2) is not equal to $\theta_{jk}$ in (1) unless either $\pi_{j}({Z})$ does not vary with ${Z}$ or $Y^{(j)}\perp{Z}$. Take the trial described in Section 2.1 and Figure 1 as an example, where $\pi_{j}({Z})$ varies with ${Z}$, the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited patients being healthier than those recruited earlier. The problem with the naive estimator $\bar{Y}_{jk}$ arises from confounding by $\pi_{j}({Z})$, not from conditioning on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$. In traditional clinical trials, $\bar{Y}_{jk}$ is valid because $\pi_{j}({Z})$ does not vary with ${Z}$.
4.2 Inverse probability weighting
As noted in Section 4.1, the naive method is biased because it fails to account for confounding by $Z$. Under Assumption 1, since the true probability of receiving treatment $\pi_{j}(Z)$ is known, this confounding can be addressed using inverse probability weighting (IPW). The identification formula is (the proof is in the Supplementary Material):
$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},\vspace{-3mm}$ (3)
where $I(\cdot)$ denotes the indicator function. Formula (3) naturally suggests the following IPW estimator of $\theta_{jk}$, which is widely used in survey sampling (Horvitz and Thompson,, 1952) and causal inference (Rosenbaum and Rubin,, 1983),
$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},\vspace{-3mm}$
where ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ and $n_{jk}$ is the number of elements in ${{{\cal I}}}_{jk}$.
This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971; Robins et al.,, 2000) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
$\hat{\theta}_{jk}^{\rm\,sipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})}.$
Both the IPW and SIPW estimators are asymptotically normal (Theorem 1 of Section 5).
The IPW and SIPW methods utilize some covariate information through weighting with $\pi_{j}({Z})$. To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2023), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect. This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (ICH E9,, 1998; EMA,, 2015; FDA,, 2021).
Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may contain or overlap with ${Z}$. We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a potentially misspecified working model. For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y_{i}$ on $\bm{X}$ using data where $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$. For binary outcomes, $\hat{\mu}_{jk}(\bm{X})$ is often obtained from a logistic working model. After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for covariate $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,aipw}=$ $\ \frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).\vspace{-3mm}$
Interestingly, in its last term, the AIPW estimator leverages covariate information from all concurrently eligible patients in ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$, including those not receiving treatments $j$ or $k$. For example, in Figure 1, when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2 using AIPW, the estimator also incorporates covariate data from subtype A patients assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2. These covariate data are not used in the IPW and SIPW estimators. Thus, AIPW offers a robust method to pool data from all concurrently eligible patients, thereby further extracting the potential of efficiency gain through platform trials.
Similarly, one can use the stabilized AIPW (SAIPW) estimator by normalizing the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$, resulting in the following SAIPW estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,saipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
Both AIPW and SAIPW estimators are asymptotically normal (see Theorem 1 in Section 5). As discussed in Section 5, whether to stabilize the weights in the AIPW approach often results in asymptotically equivalent estimators.
4.3 Post-stratification
An alternative method to account for the varying $\pi_{j}({Z})$ is to stratify subjects based on the values that $\pi_{j}({Z})$ can take. This approach is known as post-stratification in sample surveys (Fuller,, 2009) or sub-classification in the causal inference literature (Rosenbaum and Rubin,, 1984). Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible patients ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$. The resulting post-stratification (PS) estimator of $\theta_{jk}$ is
$\hat{\theta}_{jk}^{\rm\,ps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)Y_{i}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})\bar{Y}_{j}({\cal I}_{jk}^{(h)}),$
where $n({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$. This estimator does not need stabilizing, as the sum of weights is
$\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})=1.$
Compared with the SIPW, the PS requires an extra step of constructing strata for every pair $j$ and $k$.
The idea of PS also appeared in Marschner and Schou, (2022), but their approach of stratifying by enrollment window is inadequate unless ${Z}$ consists solely of enrollment time. For instance, in Figure 1, ${Z}$ includes both the disease subtype and the enrollment window; in the SIMPLIFY example in Figure 2, $Z$ is the indicator of HS and/or DA and there is only a single enrollment window. On the other hand, stratifying according to all joint levels of enrollment window and disease subtype is unnecessary. Our proposed method of stratification ensures that $\pi_{j}({Z})$ and $\pi_{k}({Z})$ remain constant within each post-stratum, which is exactly what is required to handle the issue of varying $\pi_{j}({Z})$ and $\pi_{k}({Z})$. It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section 7.1.
The PS approach utilizes covariate information by stratifying through ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$. To adjust for additional covariates, we apply AIPW separately within each stratum and then combine the results using weighted averages. This gives the following adjusted post-stratification (APS) estimator of $\theta_{jk}$,
$\hat{\theta}_{jk}^{\rm\,aps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{\cal I}}_{jk}^{(h)}}I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}+\frac{1}{n_{jk}}\sum_{i\in{{\cal I}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
Asymptotically, the PS and APS estimators are shown to be equivalent to certain types of SAIPW estimators, as detailed in Section 5.
4.4 Summary of estimators
[…]
5 Asymptotic Theory
5.1 Asymptotic normality
To develop the asymptotic theory for estimators of $\bm{\vartheta}_{jk}$ involving covariate adjustment, we need the following standard condition that the estimated working model $\hat{\mu}_{jk}$ converges to a well-defined limit $\mu_{jk}$, where the working model can be misspecified.
Assumption 2 (Stability).
For any given $j$ and $k$, there exists a function $\mu_{jk}(\cdot)$ with finite $E_{X}\{\mu_{jk}^{2}(\bm{X})\}$ such that, as $n\to\infty$, $E_{X}\{\hat{\mu}_{jk}(\bm{X})-\mu_{jk}(\bm{X})\}^{2}\to 0$ in probability with respect to the randomness of $\hat{\mu}_{jk}(\cdot)$ as a function of data, where $E_{X}$ is the expectation with respect to $\bm{X}$. If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in S1 of the Supplementary Material.
Theorem 1 shows that all six estimators described in Section 4 are consistent in estimating ${\bm{\vartheta}}_{jk}$ and are asymptotically normal under minimal statistical assumptions. In what follows, $E_{jk}=E_{kj}$, ${\rm Var}_{jk}={\rm Var}_{kj}$, and ${\rm Cov}_{jk}={\rm Cov}_{kj}$ denote respectively the expectation, variance, and covariance conditioned on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
Theorem 1.
Under Assumption 1, and Assumption 2 if adjusting for covariates, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$ whose explicit expression is given in each of the following specific case.
(a) For the
IPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ipw}$,
$\bm{\Sigma}_{jk}^{\rm\,ipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{(Y^{(j)})^{2}}{\pi_{j}({Z})},\,\frac{(Y^{(k)})^{2}}{\pi_{k}({Z})}}\right\}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T},\vspace{-2mm}$
[…]
(b) For the
SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$,
$\bm{\Sigma}_{jk}^{\rm sipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{(Y^{(j)}-\theta_{jk})^{2}}{\pi_{j}({Z})},\,\frac{(Y^{(k)}-\theta_{kj})^{2}}{\pi_{k}({Z})}}\right\}\Big{]}.\vspace{-2mm}$
(c) For the AIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,aipw}$,
$\bm{\Sigma}_{jk}^{\rm aipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{Y^{(j)}-\mu_{jk}(\bm{X})\}^{2}}{\pi_{j}({Z})},\,\frac{\{Y^{(k)}-\mu_{kj}(\bm{X})\}^{2}}{\pi_{k}({Z})}}\right\}\Big{]}+\bm{\Lambda}_{jk}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T},\vspace{-2mm}$
where $\bm{\delta}_{jk}=(\delta_{jk},\delta_{kj})^{T}$, $\delta_{jk}=E_{jk}\{Y^{(j)}\!-\mu_{jk}(\bm{X})\}$, $\mu_{jk}$ is the limit of $\hat{\mu}_{jk}$ given in Assumption 2, and $\bm{\Lambda}_{jk}$ is a symmetric matrix of order 2 with two diagonal elements $\lambda_{jk}=2\,{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\}$ and $\lambda_{kj}$, and the off-diagonal element $c_{jk}=c_{kj}=$ ${\rm Cov}_{jk}\{Y^{(j)},\mu_{kj}(\bm{X})\}+{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\}$.
(d) For the
SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$,
$\bm{\Sigma}_{jk}^{\rm saipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{Y^{(j)}-\mu_{jk}(\bm{X})-\delta_{jk}\}^{2}}{\pi_{j}({Z})},\,\frac{\{Y^{(k)}-\mu_{kj}(\bm{X})-\delta_{kj}\}^{2}}{\pi_{k}({Z})}}\right\}\Big{]}+\bm{\Lambda}_{jk}.$
(e) For the PS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$,
$\bm{\Sigma}^{\rm ps}_{jk}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}(Y^{(j)}\mid S)}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}(Y^{(k)}\mid S)}{\pi_{k}(S)}}\right\}\Big{]}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-2mm}$
where $S$ is the covariate for post-stratification, i.e., for patient $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplementary Material.
(f) For the
APS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$,
$\bm{\Sigma}^{\rm aps}_{jk}$ $=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\mid S\}}{\pi_{k}(S)}}\right\}\Big{]}$
$\quad+E_{jk}\{\bm{\Lambda}_{jk}(S)\}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-4mm}$
where $\bm{\Lambda}_{jk}(S)$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ replaced by $\lambda_{jk}(S)=2{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\mid S\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}$ and $c_{jk}(S)={\rm Cov}_{jk}\{Y^{(j)},$ $\mu_{kj}(\bm{X})\mid S\}$ $+\,{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\mid S\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}$, respectively.
All technical proofs are given in the Supplementary Material. The results in Theorem 1 for estimators with covariate adjustment hold regardless of whether the working model is correct or not. The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators in Section 6.
5.2 Efficiency comparison
As the six estimators proposed in Section 4 have different asymptotic covariance matrices, we compare the asymptotic relative efficiency between pairs of these estimators in the subsequent corollaries. The first two results focus on comparing the efficiency of stabilized versus unstabilized estimators.
Corollary 1 (Comparison of IPW and SIPW).
Under Assumption 1, the difference in asymptotic covariance matrices of the IPW and SIPW estimators is equal to
$\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$
$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.$
Since the matrix $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ in Corolary 3 is positive semidefinite (shown in S6 of the Supplementary Material) but ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ may have arbitrary signs, the asymptotic relative efficiency of IPW versus SIPW is indefinite. When ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ are both non-negative, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ is positive semidefinite and, hence, SIPW always has equal or better asymptotic efficiency than IPW in estimating any differentiable function of $\bm{\vartheta}_{jk}$. On the other hand, an example of estimating $\theta_{jk}-\theta_{kj}$ is shown in S6 of the Supplement Material in which IPW is asymptotically more efficient than SIPW.
Nonetheless, SIPW is often more efficient in simulations and real data applications (see Section 7). SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change the SIPW estimator $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect the IPW analogue $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023). For these reasons, we do not consider IPW in the remaining efficiency comparisons.
Corollary 2 (Comparison of AIPW and SAIPW).
Under Assumptions 1-2, the difference in asymptotic covariance matrices of the AIPW and SAIPW estimators is equal to
$\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}$
$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.\vspace{-2mm}$
Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition
$\delta_{jk}=\delta_{kj}=0,\vspace{-3mm}\vspace{-2mm}$ (4)
which is known as the prediction unbiasedness condition (Guo and Basse,, 2021) at the population level, where $\delta_{jk}=E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\}$ and $\delta_{kj}=E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\}$. Condition (4) holds when the working models used for covariate adjustment are correctly specified. It also holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified. Later discussions about adjustment (8) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet condition (4), it can be readily modified by (8) to satisfy this requirement. Therefore, AIPW and SAIPW are usually asymptotically equivalent and we do not include AIPW in the rest of the efficiency comparisons.
The next two results show when covariate adjustment leads to an efficiency gain. Corollary 3 compares the two stabilized IPW estimators, with or without covariate adjustment.
Corollary 3 (Comparison of SIPW and SAIPW).
Under Assumptions 1-2 and
$E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}=\delta_{jk},\ \ E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\mid S\}=\delta_{kj},$ (5)
${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{jk}(\bm{X})\mid S\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{kj}(\bm{X})\mid S\}=0,$ (6)
${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{kj}(\bm{X})\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{jk}(\bm{X})\}=0,$ (7)
where $S$ is the covariate for post-stratification as defined in Theorem 1(e), the SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ is asymptotically more efficient than the SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$, i.e.,
$\bm{\Sigma}_{jk}^{\rm\,sipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{\{\mu_{jk}(\bm{X})-\mu_{jk}\}^{2}}{\pi_{j}(S)},\,\frac{\{\mu_{kj}(\bm{X})-\mu_{kj}\}^{2}}{\pi_{k}(S)}}\right\}\Big{]}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\}$
is positive definite, where $\mu_{jk}=E_{jk}\{\mu_{jk}(\bm{X})\}$ and $\mu_{kj}=E_{jk}\{\mu_{kj}(\bm{X})\}$, unless either one of ${\rm Var}_{jk}\{\mu_{jk}(\bm{X})\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,sipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ is positive semidefinite.
When covariate adjustment for $\bm{X}$ is applied, the SAIPW estimator is not necessarily more efficient than the SIPW estimator if the working models are incorrect. Conditions (5)-(7) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the platform trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent. Actually, condition (5) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
$\tilde{\mu}_{jk}(\bm{X})=\hat{\mu}_{jk}(\bm{X})+\frac{1}{n_{jk}(S)}\sum_{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S}\{Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i})\},$ (8)
where $n_{jk}(S)$ is the number of elements in $\{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S\}$. The corresponding limit of $\tilde{\mu}_{jk}(\bm{X})$ in Assumption 2 is $\mu_{jk}(\bm{X})+E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\mid S\}$ satisfying (5) with $\delta_{jk}=\delta_{kj}=0$. Conditions (6)-(7) hold when working models are correct. Otherwise, they require certain construction of $\hat{\mu}_{jk}(\bm{X})$ and $\hat{\mu}_{kj}(\bm{X})$ that has (6)-(7) embedded. For example, (5)-(7) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2023); see S10 in the Supplement Material for details.
In fact, (5)-(7) always hold for the SAIPW estimator with $\bm{X}=S$, i.e., the SAIPW estimator with $\hat{\mu}_{jk}(\bm{X})=\bar{Y}_{j}(S)$, which is the sample mean of $Y_{i}$’s for $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}=S$. This is because, with $S$ as the stratum indicator, the limit of $\bar{Y}_{j}(S)$ is $E_{jk}(Y^{(j)}\mid\bm{X}=S)$, which is always a correct working model, and consequently conditions (5)-(7) hold. Therefore, the SAIPW estimator with $\bm{X}=S$ is asymptotically more efficient than the SIPW estimator according to Corollary 3 and it is used in later corollaries to bridge the comparison between IPW-type and PS-type estimators.
Corollary 4 compares the two post-stratification estimators, with or without additional adjustment for covariates.
Corollary 4 (Comparison of PS and APS).
Under Assumptions 1-2 and (5)-(7), the APS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,apw}$ is asymptotically more efficient than the PS estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$, i.e.,
$\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}\!=\!E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}}{\pi_{k}(S)}}\right\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}\Big{]}$
is positive definite, unless either one of ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ conditioned on $S$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}$ is positive semidefinite.
Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in the trivial scenario where either no covariate adjustment beyond $S$ in one of the treatment arms or $\pi_{j}(S)+\pi_{k}(S)=1$ and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent conditioned on $S$.
To compare the PS estimator with the IPW-type estimator, intuitively the PS estimator is asymptotically more efficient than the SIPW estimator because the former utilizes $S$ as a covariate in post-stratification whereas the latter only partially uses covariates through weighting. For a fair comparison, the PS estimator should be compared with the SAIPW estimator with $\bm{X}=S$, i.e., both estimators fully adjust for the discrete covariate $S$. This leads to the following result.
Corollary 5 (Comparison of PS and SAIPW with $\bm{X}=S$).
Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,ps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$ when $\bm{X}=S$.
Corollary 5 states that the PS estimator that stratifies on $S$ is asymptotically equivalent to the SAIPW estimator that adjusts for $S$ (as well as the corresponding AIPW estimator due to its equivalence with the SAIPW when (4) holds).
The next corollary compares APS and SAIPW with the same covariate adjustment. It demonstrates that when both estimators adjust for $\bm{X}$ using a working model that satisfies condition (5), the APS estimator is asymptotically equivalent to the SAIPW estimator.
Corollary 6 (Comparison of APS and SAIPW).
Under Assumptions 1-2 and condition (5), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,aps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$.
The last corollary completes the comparison around SAIPW with $\bm{X}=S$ and shows that under (5)-(7), adjusting for more covariates beyond just $S$ leads to additional asymptotic efficiency gain.
Corollary 7 (Comparison of SAIPW and SAIPW with $\bm{X}=S$).
Under Assumptions 1-2 and conditions (5)-(7), SAIPW with $S\subset\bm{X}$ is more efficient than SAIPW with $\bm{X}=S$.
All the efficiency comparisons are summarized in Figure 3.
Figure 3: Efficiency comparisons among all robust estimators.
6 Robust Variance Estimation
To robustly assess variability and conduct large sample inference, it is essential to obtain consistent estimators of the asymptotic covariance matrices in Theorem 1 for various estimators of $\bm{\vartheta}_{jk}$. These estimators should remain consistent whether the working models are correct or misspecified.
For IPW-type estimators, we propose the following estimators of covariance matrices:
$\hat{\bm{\Sigma}}_{jk}^{\rm\,ipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{\cal I}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)Y_{i}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)Y_{i}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}-\hat{\bm{\vartheta}}_{jk}\hat{\bm{\vartheta}}_{jk}^{T},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,sipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)(Y_{i}-\hat{\theta}_{jk})^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)(Y_{i}-\hat{\theta}_{kj})^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk}-\hat{\bm{\delta}}_{jk}\hat{\bm{\delta}}_{jk}^{T},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,saipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})-\hat{\delta}_{jk}\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})-\hat{\delta}_{kj}\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk},$
where $\hat{\bm{\vartheta}}_{jk}=(\hat{\theta}_{jk},\hat{\theta}_{kj})^{T}$, $\hat{\theta}_{jk}$ is any estimator of $\theta_{jk}$, $\hat{\bm{\delta}}_{jk}=(\hat{\delta}_{jk},\hat{\delta}_{kj})^{T}$, $\hat{\delta}_{jk}=n_{jk}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\!I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}/\pi_{j}({Z}_{i})$, $\hat{\bm{\Lambda}}_{jk}$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ respectively estimated by $\hat{\lambda}_{jk}=2\hat{q}_{jk}^{(j)}+\hat{\sigma}^{2}_{jk}$ and $\hat{c}_{jk}=\hat{q}_{kj}^{(j)}+\hat{q}_{jk}^{(k)}+\hat{q}_{jk}$, $\hat{q}_{jk}^{(j)}$ is the sample covariance of $(Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{jk}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$ and $A_{i}=j$, $\hat{q}_{kj}^{(j)}$ is $\hat{q}_{jk}^{(j)}$ with $\hat{\mu}_{jk}(\bm{X}_{i})$ switched to $\hat{\mu}_{kj}(\bm{X}_{i})$, $\hat{q}_{jk}^{(k)}$ is $\hat{q}_{jk}^{(j)}$ with $Y_{i}^{(j)}$ switched to $Y_{i}^{(k)}$, $\hat{\sigma}^{2}_{jk}$ and $\hat{\sigma}^{2}_{kj}$ are the diagonal entries and $\hat{q}_{jk}$ is the off-diagonal entry of the sample covariance matrix of $(\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{kj}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$.
For the PS and APS estimators, we propose the following estimators of $\bm{\Sigma}_{jk}^{\rm\,ps}$ and $\bm{\Sigma}_{jk}^{\rm\,aps}$,
$\hat{\bm{\Sigma}}_{jk}^{\rm\,ps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,{\rm diag}\!\left\{{\textstyle\frac{\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\sigma}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Gamma}}_{jk}$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,aps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,\Big{[}{\rm diag}\!\left\{{\textstyle\frac{\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\tau}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})\Big{]}+\hat{\bm{\Gamma}}_{jk},$
where $\hat{\pi}_{j}({\cal I}_{jk}^{(h)})=n_{j}({\cal I}_{jk}^{(h)})/n({\cal I}_{jk}^{(h)})$, $\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $Y_{i}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ estimated respectively by $\hat{\lambda}_{jk}({\cal I}_{jk}^{(h)})=2\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})+\hat{\sigma}_{jk}^{2}({\cal I}_{jk}^{(h)})$ and $\hat{c}_{jk}({\cal I}_{jk}^{(h)})=\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}^{(k)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}({\cal I}_{jk}^{(h)})$, $\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{\sigma}^{2}_{jk}({\cal I}_{jk}^{(h)})$, and $\hat{q}_{jk}({\cal I}_{jk}^{(h)})$ are respectively the same as $\hat{q}_{jk}^{(j)}$, $\hat{q}_{kj}^{(j)}$, $\hat{\sigma}^{2}_{jk}$, and $\hat{q}_{jk}$ previously defined in $\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}$ but with the index $i$ restricted to $i\in{\cal I}_{jk}^{(h)}$, $\hat{\bm{\Gamma}}_{jk}$ is the sample covariance matrix of $(\bar{Y}_{j}(S_{i}),\bar{Y}_{k}(S_{i}))$’s for $i\in{\cal I}_{jk}$, and $\bar{Y}_{j}(S_{i})$ is the sample mean of $Y_{i}$’s with $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}={\cal I}_{jk}^{(h)}$.
7 Empirical Results
7.1 Simulation study
We conduct a simulation study based on the first three enrollment windows outlined in Figure 1 to examine the finite-sample performance of all estimators presented in Section 4.
In the simulation, the observed baseline covariate vector is $\bm{W}=(X_{c},X_{b},Z_{\rm sub},Z_{\rm win})$, where $X_{c}$ is a continuous covariate uniformly distributed over the interval $(-3,3)$, $X_{b}$ is a binary covariate with $P(X_{b}=1)=0.5$, $Z_{\rm sub}$ is a binary indicator for disease subtype with $P(Z_{\rm sub}=1)=0.8$, and $Z_{\rm win}\in\{1,2,3\}$ denotes the enrollment window and satisfies
$P(Z_{\rm win}=t\mid X_{c},X_{b},Z_{\rm sub},U)=\frac{\exp(Q_{t})}{\exp(Q_{1})+\exp(Q_{2})+\exp(Q_{3})},\quad t=1,2,3,$
$Q_{1}=0.5+X_{c}+2X_{b}-Z_{\rm sub}+U$, $Q_{2}=1+2X_{c}+X_{b}-Z_{\rm sub}+U$, $Q_{3}=-0.5+X_{c}+X_{b}+Z_{\rm sub}+U$, and the unobserved baseline variable $U$ follows a standard normal distribution and is independent of $X_{c}$, $X_{b}$, and $Z_{\rm sub}$. The potential outcomes are generated as follows: $Y^{(1)}=1+X_{c}+X_{b}+Z_{\rm sub}+U+\epsilon_{1},$ $Y^{(2)}=1+X_{c}^{2}+X_{b}+Z_{\rm sub}+U+\epsilon_{2}$, $Y^{(3)}=3+X_{c}X_{b}+Z_{\rm sub}+U+\epsilon_{3}$, $Y^{(4)}=2+X_{c}Z_{\rm sub}-X_{b}+2U+\epsilon_{4}$, where $\epsilon_{j}$’s are mutually independent, each follows a standard normal distribution and is independent of $\bm{W}$ and $U$. Note that $Y^{(j)}$’s are correlated with $X_{c}$, $X_{b}$, $Z_{\rm sub}$, $U$, as well as indirectly with the enrollment time $Z_{\rm win}$.
Table 1: Assignment probabilities to each sub-study and treatment
${Z}$ Sub-study Treatment
$Z_{\rm win}$ $Z_{\rm sub}$ 1 2 3 $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$ $\pi_{4}({Z})$
1 1 0.4 0.6 0 0.50 0.20 0.30 0
1 0 1 0 0 0.50 0.50 0 0
2 1 0.3 0.3 0.4 0.50 0.15 0.15 0.20
2 0 1 0 0 0.50 0.50 0 0
3 1 0.4 0 0.6 0.50 0.20 0 0.30
3 0 1 0 0 0.50 0.50 0 0
Patients are first randomized into one of three sub-studies (labeled 1-3) depending on their enrollment windows and disease subtypes, and then to treatments within the assigned sub-study. Sub-study 1 has treatments 1 and 2, sub-study 2 has treatments 1 and 3, and sub-study 3 has treatments 1 and 4. Within each sub-study, treatments are randomized in a 1:1 ratio. The assignment probabilities are in Table 1. For the total sample size $n$ in the platform trial, we consider 500 and 1,000. When $n=500$, the expected numbers of patients per treatment are 123, 51, and 76 in sub-studies 1-3, respectively.
We consider the following estimators of the linear contrast $\theta_{j1}-\theta_{1j}$, for $j=2,3,4.$
  1. (i)
    Four IPW-type estimators. We consider the IPW, SIPW, and SAIPW estimators described in Section 4.2. For SAIPW, $\hat{\mu}_{jk}$ is obtained from fitting a linear regression of $Y$ on $(X_{c},X_{b},Z_{\rm sub})$, using data from treatment arm $j$ in ${\cal I}_{jk}$. Since condition (4) holds in this simulation, the AIPW estimator is numerically almost the same as the SAIPW estimator and thus is omitted. We add the SAIPW estimator with $\bm{X}=S$, denoted by SAIPW($S$), to check its equivalence with the PS estimator as discussed in Corollary 5. Note that the working models are correctly specified only for treatment arm 1, as the true models for the other arms are not linear in $(X_{c},X_{b},Z_{\rm sub})$.
  2. (ii)
    Four PS-type estimators. We consider the following strata for PS and APS:
    Estimating ${\small\begin{array}[]{ll}\mbox{Estimating $\theta_{21}-\theta_{12}$:}&{\cal I}_{12}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.5\,(Z_{{\rm sub},i}=0)\},\\ &{\cal I}_{12}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.2\,(Z_{{\rm win},i}=1\ \mbox{or}\ 3,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{12}^{(3)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{31}-\theta_{13}$:}&{\cal I}_{13}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.3\,(Z_{{\rm win},i}=1,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{13}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{41}-\theta_{14}$:}&{\cal I}_{14}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.2\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{14}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.3\,(Z_{{\rm win},i}=3,Z_{{\rm sub},i}=1)\}.\end{array}}\vspace{2mm}$
    The same working models $\hat{\mu}_{jk}$ for SAIPW are used for APS. For estimating $\theta_{21}-\theta_{12}$, we additionally include the PS and APS estimators that stratify on all six joint levels of ${Z}=(Z_{\rm win},Z_{\rm sub})$, denoted as PS($Z$) and APS($Z$). For estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$, we have PS($Z$) = PS and APS($Z$) = APS because using all six joint levels of ${Z}$ results in the same two strata as above (due to $\pi_{j}(Z)=0$ for some $j=3,4$).
  3. (iii)
    The naive estimator. For comparison we include the naive estimator of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, as described in Section 4.1.
  4. (iv)
    Estimators within each sub-study. For comparison, we also evaluate traditional ANOVA (sample means) and ANCOVA estimators (using linear working models to adjust for $X_{c},X_{b}$, and $Z_{\rm sub}$) based solely on data from each individual sub-study. Although these sub-study-specific estimators estimate the treatment effect within each sub-study, which differs from our target estimand defined in (1), we can still compare their variances.
The simulation results based on $5,000$ runs are reported in Table 2 for the previously described eleven estimators in (i)-(iv). The results contain the average bias, standard deviation (SD), the average of the standard error (SE), and the coverage probability (CP) of the $95\%$ confidence interval. The true values of estimands are approximated using an independent simulated dataset with size $10^{7}$. The following is a summary of the simulation results in Table 2.
Table 2: Simulation results based on $5,000$ simulation runs
$\theta_{21}-\theta_{12}=3$ $\theta_{31}-\theta_{13}=1.145$ $\theta_{41}-\theta_{14}=-0.886$
$n$ Method Bias SD SE CP Bias SD SE CP Bias SD SE CP
500 Naive -0.231 0.320 0.316 0.874 -0.185 0.342 0.340 0.916 -0.205 0.384 0.380 0.911
IPW -0.006 0.639 0.636 0.946 0.004 0.776 0.777 0.948 -0.007 0.500 0.497 0.948
SIPW -0.003 0.341 0.336 0.941 0.005 0.347 0.341 0.943 0.001 0.389 0.381 0.942
SAIPW -0.018 0.329 0.340 0.951 0.001 0.284 0.284 0.944 -0.001 0.297 0.300 0.949
$(S)$ 0.000 0.336 0.329 0.939 0.009 0.327 0.324 0.939 0.002 0.356 0.348 0.942
PS 0.000 0.336 0.335 0.945 0.009 0.327 0.330 0.949 0.002 0.356 0.356 0.946
APS -0.013 0.329 0.339 0.952 -0.001 0.286 0.289 0.947 -0.002 0.298 0.306 0.956
PS$(Z)$ Estimate or its SE cannot
APS$(Z)$ be computed in 408 runs
ANOVA* 0.001 0.354 0.350 0.946 0.006 0.421 0.417 0.945 0.004 0.424 0.425 0.946
ANCOVA* -0.008 0.329 0.321 0.942 -0.005 0.319 0.311 0.942 0.002 0.321 0.314 0.942
1000 Naive -0.230 0.226 0.224 0.819 -0.189 0.240 0.239 0.872 -0.206 0.269 0.268 0.876
IPW -0.001 0.453 0.451 0.947 0.012 0.550 0.550 0.951 0.003 0.355 0.352 0.943
SIPW 0.000 0.243 0.239 0.945 0.004 0.246 0.243 0.944 0.001 0.272 0.270 0.948
SAIPW -0.009 0.232 0.242 0.954 0.004 0.198 0.203 0.955 0.000 0.212 0.213 0.947
$(S)$ 0.001 0.238 0.234 0.945 0.004 0.233 0.230 0.942 0.003 0.252 0.247 0.943
PS 0.001 0.238 0.236 0.948 0.004 0.233 0.232 0.944 0.003 0.252 0.250 0.947
APS -0.006 0.232 0.239 0.952 0.003 0.198 0.203 0.955 0.000 0.213 0.215 0.952
PS$(Z)$ 0.001 0.238 0.236 0.948
APS$(Z)$0.005 0.233 0.229 0.942
ANOVA* 0.003 0.251 0.248 0.947 0.007 0.295 0.294 0.944 0.002 0.301 0.299 0.948
ANCOVA* -0.003 0.233 0.228 0.947 0.005 0.222 0.220 0.947 0.001 0.227 0.222 0.945
SAIPW($S$): SAIPW with $\bm{X}=S$.
PS($Z$) and APS($Z$): post-stratification using all joint levels of $Z=(Z_{\rm win},Z_{\rm sub})$ as strata (PS = PS($Z$) and APS = APS($Z$) for estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$).
*ANOVA and ANCOVA based on data within each sub-study, estimating the within sub-study treatment effects, which are $3.054$, $1.279$, and $-0.881$ for sub-studies 1-3, respectively.
  1. 1.
    The simulation results support our asymptotic theory in Section 5. Specifically, the naive estimator is biased, while all IPW-type and PS-type estimators have negligible biases; despite the misspecification of working models, covariate adjustment improves efficiency as reflected by reductions in SD; the PS and SAIPW($S$) estimators demonstrate nearly identical performance. Additionally, SEs are close to SDs and CPs are close to the nominal level 95%.
  2. 2.
    No results are reported in Table 1 for PS($Z$) and APS($Z$) at $n=500$ because either the estimate of $\theta_{jk}$ or its SE cannot be computed in 408 out of 5,000 simulation runs. This issue arises due to insufficient data in the stratum defined by $(Z_{\rm win},Z_{\rm sub})=(3,0)$, where the average number of patients is 10. A similar issue occurs in 3 out of 5,000 runs at $n=1,000$, so the results in Table 1 are based on 4,997 runs. This problem does not affect the PS and APS estimators: our proposed stratification method combines three strata $(Z_{\rm win},Z_{\rm sub})=(1,0),(2,0),(3,0)$ into ${\cal I}_{12}^{(1)}$ because they all have constant $\pi_{1}(Z)=0.5$ and $\pi_{2}(Z)=0.5$.
  3. 3.
    Between IPW and SIPW, it is clear that SIPW is much better in terms of SD under our simulation setting. Notably, IPW can have an even larger SD than ANOVA within substudies, which usually does not occur with SIPW.
  4. 4.
    The traditional ANOVA and ANCOVA estimators analyze the platform trials as if they were separate, stand-alone studies. Without covariate adjustment, the ANOVA estimator has a much larger SD than the SIPW estimator when estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$. A similar conclusion applies to the ANCOVA estimator when compared with the SAIPW or APS estimator. The efficiency difference is not as pronounced for estimating $\theta_{21}-\theta_{12}$, although the ANOVA and ANCOVA estimators still have larger SDs than the corresponding stabilized IPW-type and PS-type estimators. The reason for this is because the traditional ANOVA and ANCOVA do not utilize data across sub-studies, for example, when $n=500$, about 123, 51, and 76 patients are used under treatment arm 1 in sub-studies 1-3, respectively, whereas about 250, 121, and 150 patients in treatment arm 1 under platform trial are shared for our proposed estimators of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, respectively.
7.2 Real data analysis
As reviewed in Section 2.2 and shown in Figure 2, the SIMPLIFY design (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study. Let $Z$ be the indicator of whether a participant receives HS or DA therapy at baseline and the enrollment window, with the assignment probabilities given in Table 3. The enrollment window (EW) is treated as a categorical variable, with EW1 representing the enrollment period before the probability change, and EW2 standing for the enrollment period after the change, as presented in Figure 2. The outcome $Y$ is the 6-week mean absolute change in ppFEV1 as described in Section 2.2.
In our analysis, the treatment effects of interest are $\theta_{21}-\theta_{12}$ and $\theta_{31}-\theta_{13}$. Specifically, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS among individuals taking HS, regardless of their DA status. This represents the treatment effect among participants with $\pi_{1}(Z)>0$ and $\pi_{2}(Z)>0$, pooled across studies (i.e., participants with $\text{HS}=1$ in Table 3). The interpretation applies similarly to $\theta_{31}-\theta_{13}$. Note that our estimand is slightly different from the estimand in Mayer-Hamblett et al., (2023), as the latter analyzed the HS and DA studies separately and included participants who re-enrolled for an additional study as noted in Section 2.2. We also excluded 10 (1.7%) participants due to missing outcomes, resulting in a sample size of $n=584$: 293 in treatment 1, 119 in treatment 2, and 172 in treatment 3.
Table 3: Assignment probabilities in the SIMPLIFY trial
${Z}$ Study Treatment
HS DA EW HS DA $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$
1 1 EW1 0.50 0.50 0.50 0.25 0.25
1 1 EW2 0.75 0.25 0.50 0.375 0.125
1 0 any 1 0 0.50 0.50 0
0 1 any 0 1 0.50 0 0.50
We consider the IPW, SIPW, SAIPW, PS, and APS estimators, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section 7.1. For SAIPW and APS, in addition to $Z$, we adjust several other baseline covariates, including two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, delta F508 heterozygous, and other or unknown).
Estimates and 95% confidence intervals are presented in Figure 4, where all values are given in %. Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 4), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA. The conclusion is consistent with those in Mayer-Hamblett et al., (2023).
Figure 4: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial
8 Summary and Recommendations
Platform trials offer the potential to enhance statistical efficiency by facilitating the sharing of data between different treatment evaluations. However, realizing this potential requires careful consideration of several challenges, such as defining meaningful estimands that reflect eligibility and concurrency, and developing robust estimation techniques that accommodate the varying probabilities of treatment assignments over time and differing patient characteristics.
In this article, we explicitly present how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients. To estimate and infer this estimand with minimal statistical assumptions, we develop methods using weighting or post-stratification to account for varying treatment assignment probabilities. We also consider model-assisted approaches for covariate adjustment. We provide explicit formulas for robust variance estimation, which have been implemented in our R package, RobinCID. Through theoretical analysis and simulation studies assessing the relative efficiency of these methods, we find that the stabilized IPW (SIPW) and post-stratification (PS) significantly enhance the efficiency of the IPW estimator. Additionally, model-assisted approaches for covariate adjustment (SAIPW or APS) generally improve efficiency while maintaining robustness against model misspecification. A detailed relative efficiency comparison is in Figure 3. Note that although SIPW is less efficient than PS asymptotically, it is still useful when there are sparse strata because SIPW may have better finite-sample performance and there may be situations where the PS estimator or its variance estimator cannot be computed.

### v4

@@
 Abstract
-A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner and can accelerate the evaluation of new treatments.
-However, the flexibility that marks the potential of platform trials also creates inferential challenges.
-Two key challenges are the precise definition of treatment effects and the robust and efficient inference on these effects.
-To address these challenges, we first define a clinically meaningful estimand that characterizes the treatment effect as a function of the expected outcomes under two given treatments among concurrently eligible patients.
-Then, we develop weighting and post-stratification methods for estimation of treatment effects with minimal assumptions.
-To fully leverage the efficiency potential of data from concurrently eligible patients, we also consider a model-assisted approach for baseline covariate adjustment to gain efficiency while maintaining robustness against model misspecification.
-We derive and compare asymptotic distributions of proposed estimators in theory and propose robust variance estimators.
-The proposed estimators are empirically evaluated in a simulation study and illustrated using the SIMPLIFY trial.
-Our methods are implemented in the R package RobinCID.
-Keywords: Covariate adjustment; Eligible and concurrent controls; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
+A platform trial is an innovative clinical trial design that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments, where patients are often assigned to different subsets of treatment arms based on individual characteristics, enrollment timing, and treatment availability.
+While offering increased flexibility, this constrained and non-uniform treatment assignment poses inferential challenges, with two fundamental ones being the precise definition of treatment effects and robust, efficient inference on these effects.
+Such challenges arise primarily because some commonly used analysis approaches may target estimands defined on populations inadvertently depending on randomization ratios or trial operation format, thereby undermining interpretability.
+This article, for the first time, presents a formal framework for constructing a clinically meaningful estimand with precise specification of the population of interest.
+Specifically, the proposed entire concurrently eligible (ECE) population not only preserves the integrity of randomized comparisons but also remains invariant to both the randomization ratio and trial operation format.
+Then, we develop weighting and post-stratification methods to estimate treatment effects under the same minimal assumptions used in traditional randomized trials.
+We also consider model-assisted covariate adjustment to fully unlock the efficiency potential of platform trials while maintaining robustness against model misspecification.
+For all proposed estimators, we derive asymptotic distributions and propose robust variance estimators and compare them in theory and through simulations.
+The SIMPLIFY trial, a master protocol assessing continuation versus discontinuation of two common therapies in cystic fibrosis, is utilized to further highlight the practical significance of this research.
+All analyses are conducted using the R package RobinCID.
+Keywords: Concurrently eligible individuals; Covariate adjustment; Estimand; Inverse probability weighting; Master protocols; Relative efficiency.
 1 Introduction
-The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Saville and Berry,, 2016; Woodcock and LaVange,, 2017).
+1.1 Background
+The traditional paradigm of conducting a separate clinical trial for every investigational treatment has become ever more expensive and challenging (Woodcock and LaVange,, 2017).
 As a result, the landscape of clinical trials has been evolving, to make clinical trials faster and more efficient while continuing to provide reliable information on safety and effectiveness (The White House Office of Science and Technology,, 2023).
-The platform trial has emerged as an innovative trial design to expedite the evaluation of potentially effective treatments (Berry et al.,, 2015; Saville and Berry,, 2016; Woodcock and LaVange,, 2017; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023).
-It is a dynamic clinical trial that uses a master protocol (i.e., one overarching protocol) to evaluate multiple treatments in an ongoing manner, where new treatments can be added to the platform as they become available, and existing ones can leave the platform based on a finding of futility or efficacy.
-Such a trial structure not only streamlines operations but also has the potential to improve statistical efficiency by facilitating the sharing of data between different treatment evaluations (Mehta et al.,, 2023).
-The effectiveness of platform trials in addressing urgent medical challenges quickly is evidenced by the hallmark RECOVERY platform trial, which was designed to rapidly evaluate COVID-19 therapies during the pandemic (Normand,, 2021; Buenconsejo et al.,, 2023).
-Platform trials are also attractive in the exploratory phase to quickly screen multiple treatments for diseases (Wang and Yee,, 2019) and in rare diseases to enhance the chance of identifying effective treatments (Dhaenens et al.,, 2024).
-Two examples of platform trials are presented in Section 2 to provide detailed illustrations.
-Although platform trials offer several advantages over traditional trials, they also create challenges in statistical inference on treatment effects, as highlighted in a recent Food and Drug Administration draft guidance (FDA,, 2023).
-One primary challenge is how to define a meaningful treatment effect in a platform trial setting, as multiple treatments may target different biomarker-defined subgroups and enter the trial at different times — a complexity that does not exist in traditional trials.
-In platform trials, when comparing a control and a treatment of interest, many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, (2023) have recommended using eligible and concurrent controls, which refer to the patients who are eligible to receive the treatment and are concurrently randomized to the control arm during the same enrollment period as the other patients being assigned to the treatment arm.
-The use of eligible and concurrent controls, or concurrently eligible patients, preserves the integrity of randomized comparisons and inference on treatment effects.
-In contrast, patients who are either ineligible or nonconcurrent may not be comparable to the treated group due to measured and unmeasured differences (FDA,, 2023).
-In Section 3, we explicitly present, for the first time in the literature, how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients, which aligns with the ICH E9 , 2019 (R1) estimand framework.
-The next challenge is on robust estimation of the defined treatment effects for concurrently eligible patients.
-Given the adaptive nature of platform trials, where treatments can be added and removed, the available treatment options vary over time.
-This dynamic, coupled with different eligibility criteria for each treatment, means that the probabilities of treatment assignments can vary over time and patient characteristics — another complexity absent in traditional trials.
-Thus, naively comparing outcome means between two treatment arms may yield biased results, an example of Simpson’s paradox (Collignon et al.,, 2020; Park et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021).
-The literature on statistical methods for platform trials is expanding, though it primarily focuses on model-based techniques for incorporating nonconcurrent controls, which can be sensitive to model misspecification (Saville et al.,, 2022; Bofill Roig et al.,, 2022; Elm et al.,, 2012; Wang et al.,, 2023).
-Some research has been conducted on statistical methods for concurrently eligible patients.
-Ventz et al., (2018) considered several adaptive randomization procedures and tests for the hypothesis of no treatment effect.
-Marschner and Schou, (2022) proposed stratification using enrollment windows, without taking into account other factors related to treatment and randomization.
-Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows.
-However, there remains a lack of rigorous investigation on how to make robust inference on treatment effects in platform trials under the same minimal statistical assumptions as those in traditional clinical trials.
-To fill this gap, in Section 4, we develop methods for estimating treatment effects for concurrently eligible patients using inverse probability weighting or post-stratification to account for varying treatment assignment probabilities.
-Additionally, we consider model-assisted approaches for covariate adjustment to gain efficiency while maintaining robustness against model misspecification.
-In Section 5, we theoretically establish the asymptotic distributions of proposed estimators and compare their asymptotic relative efficiency.
-We provide robust variance estimators in Section 6.
-Simulation results and an application of our methods to the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) are in Section 7.
-Our methods are implemented in the R package RobinCID.
-2 Examples of Platform Trials
-2.1 A stylistic example
-First, we present a stylistic example of a platform trial in Figure 1 to provide a concrete illustration of its key features.
-This trial comprises five sub-studies, each containing a randomized controlled comparison of several treatments.
+Master protocol trials have emerged as innovative designs to accelerate the evaluation of potentially effective treatments (Woodcock and LaVange,, 2017).
+These trials evaluate one or more treatments across one or more diseases within an overarching study structure.
+Three major types of master protocol designs are umbrella trials, platform trials, and basket trials.
+Umbrella trials evaluate multiple therapies for a single disease (Rashdan and Gerber,, 2016; Renfro and Sargent,, 2017), basket trials evaluate a medical product for multiple diseases or disease subtypes (Redig and Jänne,, 2015; Hobbs et al.,, 2022), while platform trials offer the added flexibility of adding or removing treatment arms over time (Berry et al.,, 2015; Saville and Berry,, 2016; Park et al.,, 2020; Gold et al.,, 2022; Burki,, 2023; Pitre et al.,, 2023).
+These three types are not mutually exclusive, and a master protocol trial can incorporate features of more than one type.
+For example, an umbrella-platform hybrid for lung cancer might test different treatments across biomarker-defined subtypes (umbrella) while allowing treatments to enter or exit over time (platform).
+Similarly, a basket-platform hybrid might study a single targeted therapy across multiple cancers (basket), with new cancer types or biomarkers added adaptively over time (platform).
+Therefore, umbrella and basket trials can be viewed as special cases of platform trials, and, in this paper, we use the terms platform trial and master protocol interchangeably.
+Master protocols originated in oncology and remain widely used in this field, where many new biomarker-driven therapies require efficient evaluation (Park et al.,, 2019, 2020; Ouma et al.,, 2022).
+By utilizing a common screening platform to efficiently screen for biomarkers and identify all treatments for which a patient is eligible, these trial designs significantly improve operational efficiency, accelerate patient accrual, and enable the study of rare tumor subtypes within one trial (Woodcock and LaVange,, 2017).
+In addition to streamlining operations, these designs also have the potential to improve statistical efficiency by facilitating data sharing across different treatment evaluations.
+Notable examples include the BATTLE umbrella trials (Kim et al.,, 2011; Papadimitrakopoulou et al.,, 2016), the Lung-MAP platform trial (Herbst et al.,, 2015), and the I-SPY 2 platform trial (Barker et al.,, 2009).
+The utility of master protocol trials is much broader.
+They are gaining prominence in other diseases areas, including infectious diseases (Normand,, 2021), neurodegenerative diseases such as amyotrophic lateral sclerosis (Quintana et al.,, 2023) and Alzheimer’s disease (Bateman et al.,, 2017), as well as inflammatory bowel disease (Honap et al.,, 2024).
+Section 2 presents two trials to provide concrete illustrations.
+1.2 Two fundamental gaps
+Compared to traditional parallel-group trials, master protocol trials often share a key feature: constrained and non-uniform treatment assignment.
+Constrained means some participants have zero probability of receiving certain treatments, while non-uniform refers to varying non-zero assignment probabilities based on participant characteristics, enrollment time, or treatment availability.
+This feature marks a key distinction between master protocol and traditional trial designs, and poses significant challenges for the interpretability of and statistical inference for the treatment effect, some of which are highlighted in a recent draft guidance from the U.S.
+Food and Drug Administration (FDA, 2023b, ).
+Although master protocols are increasingly conducted, two fundamental questions arising from their constrained and non-uniform treatment assignment remain unresolved.
+1. How to define a clinically meaningful estimand?
+An estimand is “a precise description of the treatment effect reflecting the clinical question posed by a given clinical trials objective” and is of ultimate importance (ICH E9 , 2019, R1).
+In master protocol trials, treatments often have different eligibility criteria and can be added or removed over time, leading to changes in the set of available treatment options throughout the trial.
+Consequently, the set of treatments to which an individual can be randomized and the corresponding assignment probabilities can vary over time and across individual characteristics.
+This leads to constrained and non-uniform treatment assignments.
+As a result, defining an estimand becomes more challenging because the treatment effect must be interpreted in the context of restricted eligibility, variable assignment probabilities, and evolving trial structure over time.
+In particular, among the five estimand attributes (ICH E9 , 2019 (R1)), the target population of interest requires special attention.
+Many researchers (Lee and Wason,, 2020; Dodd et al.,, 2021) and FDA, 2023b have recommended using a population of individuals who are concurrently randomized and meet the eligibility criteria, as this preserves the integrity of randomized comparisons and avoids systematic differences between treatment groups due to measured and unmeasured factors.
+However, this general recommendation does not uniquely define the population, leading to ambiguity.
+For instance, in Cases I-II in Section 3.2, a seemingly reasonable population (and thus the estimand) may depend arbitrarily on randomization probabilities, even though it technically follows the general recommendation in FDA, 2023b .
+To address this question, Section 3.2 introduces a population for comparing two treatments (typically with one as a control) that not only follows the general recommendations in FDA, 2023b but also respects the eligibility criteria while avoiding arbitrary dependencies on randomization probabilities.
+We refer to this as the entire concurrently eligible (ECE) population.
+This definition also applies to traditional parallel-group trials, to which we draw simple connections using Case I in Section 3.2 to explain why this population is sensible.
+The corresponding estimand for the ECE population is defined in Section 3.3.
+As in traditional trials, the target population in a master protocol trial may sometimes be a specific subgroup or a re-weighted population.
+In such cases, the ECE population can serve as the reference and may be restricted to the relevant subgroup or re-weighted to resemble the target population.
+Our proposal is the first in the literature to offer a systematic framework for defining clinically meaningful estimands with clear specification of the population of interest.
+Given the central importance of the estimand, our contribution enhances the fundamental understanding of master protocol trials and lays the groundwork for future statistical research—both frequentist and Bayesian—on topics such as interim analysis and the use of nonconcurrent controls.
+2. How to provide robust estimation and inference regarding the defined treatment effect?
+Although randomization is preserved, the commonly used estimator in traditional trials—comparing outcome means between treatment arms—can easily lead to biased estimates in master protocol trials, where treatment assignment is often constrained and non-uniform, potentially giving rise to issues such as Simpson’s paradox (Collignon et al.,, 2020; Dodd et al.,, 2021; Lee et al.,, 2021; Meyer et al.,, 2021).
+Consequently, estimating treatment effects in master protocol trials is not straightforward.
+A promising first step is to adapt inverse probability weighting or post-stratification to account for the constrained and non-uniform treatment assignment.
+Additionally, model-assisted covariate adjustment can offer another means to further unlock the efficiency potential of master protocol trials, while maintaining robustness against model misspecification.
+To guide practitioners in selecting among these methods, we derive the asymptotic variance and relative efficiency of these estimation methods.
+These results are novel and, to our knowledge, have not previously appeared in the literature.
+1.3 Prior work
+The increasing popularity of platform trials has led to a surge of research addressing their statistical challenges in both design and analysis.
+Many studies have focused on adaptive designs, which aim to efficiently screen candidate treatments and adaptively randomize individuals to the most promising treatments (Yuan et al.,, 2016; Kaizer et al.,, 2018; Ventz et al.,, 2018).
+Additionally, Bofill Roig et al., (2024) examined the optimal allocation ratios in these trials.
+A key concern in the analysis of platform trials is the potential presence of time trends, which can introduce bias when using nonconcurrent controls (Bofill Roig et al.,, 2023).
+To address this, various strategies have been proposed to account for time trends in the analysis (Elm et al.,, 2012; Saville et al.,, 2022; Bofill Roig et al.,, 2022; Jiang et al.,, 2023; Wang et al.,, 2023; Guo et al.,, 2024).
+However, all of these statistical methods have predominantly relied on correct modeling and are sensitive to model misspecification.
+The platform trial literature on non-model-based approaches for concurrently eligible individuals remains limited.
+Marschner and Schou, (2022) proposed stratification by enrollment windows.
+Huang et al., (2023) developed inference methods for time-to-event outcomes under a strong constancy assumption that, conditional on a discrete baseline covariate, the efficacy of treatments is invariant across enrollment windows.
+More recently, Santacatterina et al., (2025) defined an estimand based on the concurrent period and established identification and estimation results using only concurrent controls.
+They also proposed strategies to incorporate nonconcurrent controls under a strong conditional exchangeability assumption between concurrent and nonconcurrent controls, given observed covariates.
+However, none of these studies consider the more general setting in which treatments have distinct eligibility criteria, nor do they rigorously investigate how to define a clinically meaningful estimand within a general framework that is independent of specific study designs or modeling assumptions.
+1.4 Outline of the article
+Section 2 presents the SIMPLIFY trial and a stylized trial used to emphasize key issues.
+Section 3 details the ECE population and corresponding estimand.
+In Section 4, we develop robust estimation methods, without or with covariate adjustment.
+In Section 5, we establish the asymptotic distributions and relative efficiency of the proposed estimators.
+We highlight the practical significance through the SIMPLIFY trial in Section 6, and we conclude with a summary and recommendations in Section 7.
+2 Real and Stylistic Trials
+2.1 The SIMPLIFY trial
+We begin by presenting the real master protocol trial that motivates this work: the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023).
+Although it resembles an umbrella trial, where multiple treatments are evaluated concurrently and have different eligibility criteria, it also presents a key feature of platform trials: varying randomization ratio over time.
+Despite its seemingly simple design diagram, the trial involves both constrained and non-uniform treatment assignments.
+Its clarity makes it ideal for highlighting the two fundamental gaps, as it facilitates understanding of the underlying challenges.
+Cystic fibrosis (CF) is caused by mutations in the CF transmembrane conductance regulator (CFTR) gene.
+CFTR modulator therapies are a class of small-molecule drugs designed to address the underlying cause of CF, leading to transformative improvements in patient outcomes.
+A landmark example is the triple-combination therapy elexacaftor/tezacaftor/ivacaftor (ETI), approved for individuals with at least one F508del allele, a mutation present in over 85% of people with CF.
+Given the potential overlap in therapeutic effects between ETI and pre-existing treatments such as hypertonic saline (HS) and dornase alfa (DA) and the need to reduce treatment burden, it is important to evaluate the impact of discontinuing these pre-existing therapies in individuals taking ETI.
+The SIMPLIFY trial was conducted within this context.
+It includes two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the CF Therapeutics Development Network.
+The primary objective for each study is to determine whether discontinuing either HS or DA, respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator ETI.
+The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1).
+Figure 1: SIMPLIFY design schema.
+Individuals on a single therapy enter into the corresponding study with probability 1, while those on both are randomized to the HS or DA study, with ratio of 1:1 during enrollment window (EW) 1, then 3:1 during EW 2.
+Within HS (or DA) study, individuals are randomized 1:1 to continue or discontinue HS (or DA).
+Table 1: Assignment probabilities in the SIMPLIFY trial.
+Baseline variables EW Studies Treatments
+HS DA EW HS study DA study 1 2 3
+1 1 EW1 0.50 0.50 0.50 0.25 0.25
+1 1 EW2 0.75 0.25 0.50 0.375 0.125
+1 0 any 1 0 0.50 0.50 0
+0 1 any 0 1 0.50 0 0.50
+Figure 1 depicts the SIMPLIFY study design.
+All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups based on medication use when entering the trial: those on HS only, DA only, or both therapies.
+Individuals on a single therapy (HS only or DA only) are assigned to the corresponding study with probability 1.
+Those on both therapies are randomized to the HS or DA study, with 1:1 ratio during initial enrollment window 1 (before January 14, 2022), and 3:1 ratio during enrollment window 2 (after January 14, 2022) to assign more participants to the HS study, as the DA study was enrolling more quickly.
+This swift shift in the randomization ratio exemplifies how the flexibility of master protocol trials can accelerate clinical trials.
+Within each study (HS or DA), participants were then randomized 1:1 to either continue or discontinue their respective therapy.
+Table 1 summarizes assignment probabilities of the SIMPLIFY trial.
+It illustrates both constrained and non-uniform treatment assignment: participants receiving only HS or DA have zero probability of being assigned to certain treatment arms, and the non-zero probabilities of receiving any given pair of treatments may vary based on baseline characteristics and enrollment windows.
+Detailed Analysis is carried out in Section 6.
+2.2 A stylistic trial
+To motivate a general statistical framework for master protocol trials, we introduce a stylistic trial in two operation formats, both can be unified under our proposed framework.
+This unification highlights an important point: while trial design may influence clinical operations, it should not alter the clinical questions being addressed.
+Figure 2(a) presents the stylistic trial in a sub-study format, where treatments are evaluated in separate sub-studies.
+Figure 2(b) shows a multi-arm format, where each treatment is studied in a separate arm without dividing the trial into sub-studies.
+The master protocol trial (in sub-study format) in Figure 2(a) comprises four sub-studies, each containing a randomized controlled comparison of two treatments.
 Initiating or stopping a sub-study is governed by a master protocol.
 Additionally, each sub-study has an intervention-specific appendix outlining protocol elements specific to that sub-study.
-Patients are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis.
+Individuals are randomized first to eligible sub-studies and subsequently to treatment arms within a sub-study, which gives explicit treatment allocation probabilities for every subject, sufficient for the identification of treatment effects and statistical analysis.
 Real trials resembling this structure include Lung-MAP for lung cancer (Herbst et al.,, 2015; Redman et al.,, 2020) and an ongoing trial for chronic pain (ClinicalTrials.gov, NCT05986292).
-Figure 1: A stylistic example of a platform trial.
-The x-axis represents calendar time discretized into windows with different potential enrollment in components of the platform trial.
-The figure illustrates both different treatment comparisons and different patient subgroups evolving over time.
-In Figure 1, each box represents a sub-study along with its status, and each striped block within a box represents the enrollment period for a treatment evaluated in that sub-study.
-The horizontal axis denotes calendar time.
-Enrollment windows are defined by changes in the available treatments for enrollment.
-For instance, the initiation of enrollment for sub-study 3 marks the start of enrollment window 2, within which treatment 4 is compared to treatment 1 as the control among patients with disease subtype A.
-Likewise, the conclusion of the enrollment period for sub-study 2 defines the start of enrollment period 3.
-Subsequently, the start of treatment 3 as the new control in place of treatment 1 among subtype A defines the start of enrollment window 4.
-The following is a narrative of Figure 1 according to the order of enrollment windows.
-Enrollment Window 1.
-In this example, there are two initial sub-studies.
-Patients with any disease subtype is eligible to sub-study 1, while sub-study 2 is restricted to patients with subtype A.
-During enrollment window 1, patients with subtype A are randomly assigned to sub-study 1 or 2, and then randomized to treatment 1 or 2 if they are in sub-study 1 and to treatment 1 or 3 if they are in sub-study 2.
-Patients with disease subtypes other than A are enrolled in sub-study 1 and then randomized to treatment 1 or 2.
-Enrollment Window 2.
-Sub-study 3 is added to the platform trial when a new intervention treatment 4 becomes available for disease subtype A.
-Following this addition, a patient with subtype A is randomly assigned to either sub-study 1, 2, or 3, and subsequently randomized to the treatments being assessed within the respective sub-study.
-Enrollment Window 3. After sub-study 2 reaches full enrollment, it closes its enrollment. Later during this period, treatment 3 meets the pre-specified success criteria.
-Enrollment Window 4.
-Since treatment 3 has been demonstrated to be superior to treatment 1 for subtype A, it replaces treatment 1 as the control for subtype A.
-Consequently, treatment 3 replaces treatment 1 in sub-study 3.
-Moreover, sub-study 1 excludes patients with subtype A, and sub-study 4 is introduced to continue evaluating treatment 2 in comparison to the new control treatment 3 for patients with subtype A.
-Enrollment Window 5.
-Sub-study 5 is added upon the introduction of a new treatment 5 for patients positive for biomarker B.
-In sub-study 5, treatment 5 is evaluated against treatment 1 for patients without subtype A, or to treatment 3 for patients with subtype A.
-Enrollment Window 6. Sub-study 4 has finished its planned enrollment and is no longer accepting new patients.
-Enrollment Window 7.
-A new treatment, treatment 6, has been introduced in sub-study 5 for patients positive for biomarker B.
-During this enrollment window, treatment 2 in sub-study 4 fails to meet the pre-specified success criteria.
-As the platform trial continues, additional sub-studies or treatments may be added as new interventions emerge for one or more targeted groups of patients, while existing sub-studies may be concluded due to either meeting success criteria or demonstrating futility.
-As exemplified in Figure 1, platform trials can be very flexible: adding new sub-studies and/or treatments with different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial.
-Moreover, compared to a traditional clinical trial, it enables potential pooling of data across sub-studies.
-Specifically, in traditional clinical trials, patients self-select into studies and are randomized into treatment arms, with treatment comparisons conducted solely within each study.
-In contrast, platform trials enroll patients into a master protocol, randomizing them into eligible sub-studies and subsequently treatment arms, and consequently, patients can be pooled across sub-studies.
-For example, when comparing treatments 1 and 2, in a platform trial like Figure 1, patients from not only sub-study 1 but also other sub-studies receiving treatment 1 can be included in the analysis, whereas we analyze data within sub-study 1 in a traditional trial.
-See the simulation results and discussions in Section 7.1 for the efficiency gain.
-2.2 A real example: the SIMPLIFY trial
-Our second example is a real platform design aliased SIMPLIFY for people with cystic fibrosis (Mayer-Hamblett et al.,, 2023).
-It contains two concurrently run multi-center randomized controlled studies at 80 participating clinics across the United States in the Cystic Fibrosis Therapeutics Development Network.
-The primary objective for each study is to determine whether discontinuing either hypertonic saline (HS) or dornase alfa (DA), respectively, is non-inferior to continuing therapy in individuals using the highly effective modulator, elexacaftor plus tezacaftor plus ivacaftor (ETI).
-The outcome is the 6-week mean absolute change in percent-predicted forced expiratory volume in one second (ppFEV1).
-Figure 2 depicts the SIMPLIFY study design.
-Figure 2: SIMPLIFY study design.
-All enrolled individuals have initiated ETI for at least 90 days and belong to one of three groups: those receiving HS only, those receiving DA only, and those receiving both therapies.
-Individuals on a single therapy enter the corresponding study, while those on both are randomized to the HS or DA study, first with a 1:1 probability and then with a 3:1 probability during subsequent enrollment.
-All individuals in HS study (or DA study) are randomized 1:1 to continue or discontinue HS (or DA).
-3 Setup and Estimand
-3.1 Randomization procedures in platform trials
-Consider a platform trial with $J$ treatments and $n$ patients.
-For each patient $i$, let $\bm{W}_{i}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A_{i}$ be the treatment assignment indicator that equals $j$ if the patient is assigned to arm $j$, $Y_{i}^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y_{i}=Y_{i}^{(A_{i})}$ be the observed outcome.
-Throughout, $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, are independent and identically distributed realizations of $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments.
-The following assumption describes the randomization procedure in platform trials.
-Assumption 1 (Randomization).
-There exists an observed baseline discrete variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
-Assumption 1 highlights a fundamental difference between the randomization procedure used in platform trials versus traditional clinical trials.
-In traditional trials, the randomization probability is typically a fixed positive constant.
-However, in platform trials, this probability, $\pi_{j}(Z)$, varies depending on a baseline discrete variable $Z$ and may be zero for some treatments.
-Notably, $Z$ usually includes the enrollment window by calendar time and some key patient characteristics, as the enrollment window determines which treatments are available for enrollment, and patient characteristics determine the treatments for which a patient is eligible.
-Hence, $\pi_{j}(Z)$ may be zero if treatment $j$ has not entered the platform trial or has closed its enrollment, or if the patient does not meet the eligibility requirements for treatment $j$.
-Additionally, non-zero values of $\pi_{j}(Z)$ can differ over time and patient characteristics.
-Values of $\pi_{j}({Z})$, $j=1,...,J$, are known by design, but $\pi_{j}({Z})$ should be viewed as random due to the randomness of ${Z}$.
-The example in Figure 1 illustrates these dynamics.
-Here, $Z=(Z_{\rm win},Z_{\rm sub},Z_{\rm B+})$ represent the enrollment window by calendar time, disease subtype (A or not A), and biomarker B (positive or negative), respectively, where $Z_{\rm B+}$ does not affect treatment assignment until enrollment window 5.
-For simplicity, consider randomization with equal probabilities.
-If patient $i$ with subtype A enters the trial during enrollment window 1 in which only sub-studies 1 and 2 are open, this patient is randomized to either sub-study with equal probability.
-If randomized to sub-study 1, the patient is further randomized to treatment 1 or 2 with equal probability.
-If randomized to sub-study 2, the patient is further randomized to treatment 1 or 3 with equal probability.
-Consequently, for this patient $i$, $\pi_{1}({Z}_{i})=0.5\times 0.5+0.5\times 0.5=0.5$, $\pi_{2}({Z}_{i})=0.5\times 0.5=0.25$, $\pi_{3}({Z}_{i})=0.5\times 0.5=0.25$, with all other $\pi_{j}({Z}_{i})$ values equal to 0.
-If another patient $i^{\prime}$ without subtype A enters the trial during the same enrollment window 1, then this patient is assigned to sub-study 1 with probability 1.
-Consequently, $\pi_{1}({Z}_{i^{\prime}})=\pi_{2}({Z}_{i^{\prime}})=0.5$ and all other $\pi_{j}({Z}_{i^{\prime}})$ values are 0.
-Similarly, if another patient $i{{}^{\prime\prime}}$ with subtype A enters the trial during enrollment window 2 in which sub-studies 1,2,3 are open, this patient has $\pi_{1}(Z_{i^{\prime\prime}})=0.5$, $\pi_{2}(Z_{i^{\prime\prime}})=\pi_{3}(Z_{i^{\prime\prime}})=\pi_{4}(Z_{i^{\prime\prime}})=1/6$, with all other $\pi_{j}({Z}_{i^{\prime\prime}})$ values equal to 0.
-These cases exemplify how the positivity and values of $\pi_{j}(Z)$ can vary with $Z$.
-3.2 Treatment effects for concurrently eligible patients
-We first define a clinically meaningful estimand (ICH E9 , 2019, R1) for comparing a given pair of treatments, $j$ and $k$, one of which is typically a control.
-As discussed in the introduction, many researchers and the FDA have recommended that the comparison be based on concurrently eligible patients, i.e., patients who meet the eligibility criteria for both treatments and are enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment $j$ or $k$.
-Formally, the concurrently eligible patients for treatments $j$ and $k$ are those with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
-We provide two examples to illustrate this concept.
-The first example is based on Figure 1.
-When an investigator compares treatments 1 and 3 at the end of sub-study 2, the concurrently eligible patients are all subtype A patients from enrollment windows 1 and 2, which includes all patients in sub-study 2, as well as patients with subtype A in sub-study 1 and all patients in sub-study 3 from the first two windows.
-Patients without subtype A or enrolled in later windows are not included, as they are either ineligible or nonconcurrent.
-The second example is based on the SIMPLIFY design in Figure 2.
-The original analysis by Mayer-Hamblett et al., (2023) calculated the effects of discontinuing vs.
-continuing HS or DA within each study separately.
-However, this approach did not utilize all concurrently eligible patients with additional control data that could have been used from participants assigned to continue therapy across either study.
-Thus, when comparing discontinuing HS versus continuing therapy, for instance, the concurrently eligible patients are all individuals on HS regardless of their DA status, which includes all participants in the HS study, as well as those on both HS and DA in the DA study.
-We propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ among concurrently eligible patients:
+A narrative of Figure 2(a) is included in Section S1 of the Supplement.
+Figure 2: A stylistic master protocol trial in two operating formats: (a) sub-study format, (b) multi-arm format.
+Each box in (a) represents a sub-study, along with its eligibility criteria and status.
+Each striped block within a box in (a) or standalone in (b) represents the enrollment period for a treatment evaluated.
+The horizontal axis shows calendar time, divided into discrete windows based on changes in available treatments for enrollment.
+In Figure 2(b), we reframe the same master protocol trial into a multi-arm format.
+In enrollment window 1, there are three initial treatment arms: investigational treatments 2 and 3 and treatment 1 serving as the common control.
+In this window, individuals with subtype A can be randomized to any of the three arms, while those with subtypes other than A can only be randomized to treatments 1 or 2.
+Descriptions for subsequent enrollment windows are omitted due to their similarity with Figure 2(a).
+Examples of real trials that follow this structure include the STAMPEDE trial for prostate cancer (Sydes et al.,, 2012) and the RECOVERY trial for COVID-19 therapies (Normand,, 2021).
+As exemplified in Figure 2, master protocol trials can be very flexible: adding new sub-studies and/or treatments with new interventions and different eligibility, dropping treatments or sub-studies, and changing the control in the middle of the trial.
+3 Estimand
+3.1 Randomization procedure
+Consider a platform trial with $J$ treatments.
+For each individual, let $\bm{W}$ be a vector of all observed baseline variables (e.g., baseline covariates and enrollment window by calendar time) not affected by the treatment, $A$ be the treatment assignment indicator that equals $j$ if the individual is assigned to arm $j$, $Y^{(j)}$ be the potential outcome under treatment $j$, $j=1,...,J$ (Neyman,, 1923; Rubin,, 1974), and $Y=Y^{(A)}$ be the observed outcome.
+We assume throughout that $(\bm{W}_{i},Y_{i}^{(1)},\dots,Y_{i}^{(J)},A_{i})$, $i=1,\dots,n$, is an independent and identically distributed sample of size $n$ from $(\bm{W},Y^{(1)},\dots,Y^{(J)},A)$ with finite second-order moments.
+The following is assumed for randomization in platform trials.
+Assumption 1 (Constrained and non-uniform randomization).
+There exists an observed baseline variable ${Z}$ such that $A\perp(\bm{W},Y^{(1)},\dots,Y^{(J)})\mid{Z}$ and $P(A=j\mid{Z})=\pi_{j}({Z}),$ where $\perp$ stands for independence, $\mid$ stands for conditioning, $10\leq\pi_{j}({Z})<$ 1, $j=1,...,J$, are known treatment assignment probabilities once ${Z}$ is known, and $\sum_{j=1}^{J}\pi_{j}({Z})=1$.
+Assumption 1 highlights a fundamental difference between platform and traditional trials.
+Traditional trials typically use a fixed randomization probability, whereas in platform trials, the randomization probabilities, $\pi_{j}(Z)$, vary with a baseline variable $Z$, which usually is discrete and may include enrollment window (defined by calendar time), site, and individual characteristics.
+$\pi_{j}(Z)$ may even be zero for some treatments.
+For example, if treatment $j$ is not yet available, has closed enrollment, is not offered at a given site, or the individual is ineligible.
+Even when non-zero, $\pi_{j}(Z)$ can vary across individuals and over time.
+Although $\pi_{j}({Z})$ is known by design, it is random due to the randomness of ${Z}$.
+[…]
+3.2 Population
+Building on the discussion in Section 1.2, we now formally define the population of interest for comparing two treatments $j$ and $k$ (one of which is typically a control).
+In this paper, population specifically refers to the population of individuals used to define the estimand.
+Definition (ECE population).
+For the given treatments $j$ and $k$, the entire concurrently eligible (ECE) population is a population of all individuals who meet the eligibility criteria for both treatments and could potentially be enrolled during a time period when both treatments are available, and therefore could have been randomized to either treatment.
+Formally, the ECE population is a population of all individuals with $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
+The ECE population is a population of all individuals who could potentially be assigned to treatments $j$ and $k$, rather than the population represented by those actually assigned to these treatments (i.e., the population represented by $A\in\{j,k\}$).
+In traditional trials, this distinction is often overlooked because the potential and actual populations coincide.
+However, in platform trials, this issue has been brought to the forefront.
+We present two basic cases to illustrate why using the population represented by those actually assigned to the two treatments or sub-studies may lead to unexpected problems.
+In Case I, all participants are eligible for all treatments, and the focus is on how changing treatment allocation probabilities over time (non-uniform treatment assignment) can cause issues in defining the population.
+In Case II, all treatments enroll concurrently, and the focus is on how different eligibility criteria for different treatments (constrained treatment assignment) complicates population definition.
+Finally, we revisit the SIMPLIFY trial, which features both constrained and non-uniform treatment assignment, reflected in time-varying treatment allocation probabilities and varying eligibility criteria across treatments.
+Figure 3: Comparison of traditional parallel-group and platform trial designs in a hypothetical mild to moderate disease population (Case I).
+In this case, the ECE population for comparing treatments 2 and 1 coincides with the entire trial population across all designs.
+In contrast, in platform trials, the population represented by those assigned to specific treatments or sub-studies may vary arbitrarily depending on the trial design or randomization probabilities.
+Case I in Figure 3 shows various trial designs evaluating treatments 1 and 2, with all participants eligible for all treatments.
+The entire trial population is 50% mild and 50% moderate, but patient characteristics shift over time: enrollment window 1 representing one third of the population consists of 30% mild and 70% moderate, while enrollment window 2 comprising the remaining two third includes 60% mild and 40% moderate.
+We focus on the population for comparing treatments 2 and 1, as these two treatments are open for enrollment during the same period across all designs.
+If a traditional parallel-group trial design (Design 1) is used, all participants have a positive probability of receiving either treatment, making the ECE population identical to the entire trial population (50% mild and 50% moderate).
+This is typically the default population for comparing treatments 1 and 2.
+The same applies to all platform trial designs in Figure 3, where the ECE population remains identical to the entire trial population regardless of the specific platform trial type (Design 2 vs.
+3) or randomization probabilities (Design 2 vs.
+2’).
+This consistency provides a principled justification for using the ECE population, as the target population for treatment comparisons should not depend on the specific trial operation formate or randomization probabilities, in the same way that choices of randomization ratio do not affect the study population in traditional trials.
+However, in platform trials, it is common to conduct analyses that naively subset participants assigned to treatment 1 or 2 (Design 2 and 2’), or to sub-study 1 (Design 3).
+Even without explicit subsetting, methods such as the t-test are often applied in Designs 2 and 2’, where the randomization ratio between treatments 1 and 2 is constant.
+These analyses target the population of individuals assigned to specific treatments or sub-studies.
+Although such analysis populations still consist of concurrently eligible individuals, as shown in Figure 3, they differ from the ECE population and can be unexpectedly influenced by design formats and randomization ratios.
+See Section S1 of the Supplement for more discussions.
+Lastly, we remark that the practice of focusing on the analysis populations assigned to specific treatments or sub-studies likely originates from traditional parallel-group trials.
+In traditional trials, these issues do not arise because the constant randomization ratio ensures that the analysis population in any arm coincides with the ECE population.
+Figure 4: A hypothetical umbrella trial (Case II) with a population with different disease severity (mild, moderate, and severe).
+In this case, the ECE population for comparing treatments 2 and 1 is the mild-to-moderate disease population (with 50% mild and 50% moderate).
+In contrast, the analysis population defined by participants assigned to sub-study 1 may vary arbitrarily depending on how individuals with moderate disease are randomized between the two sub-studies.
+Case II considers an umbrella trial where all treatments enroll concurrently.
+There are two sub-studies: one evaluates treatments 1 and 2 in a population with mild to moderate disease, and the other evaluates treatments 3 and 4 in a population with moderate to severe disease.
+The moderate disease population is randomized between the two sub-studies with a 1:1 ratio.
+These two populations are commonly of interest in the context of ulcerative colitis (FDA,, 2022).
+Suppose the overall disease population consists of 40% mild, 40% moderate, and 20% severe disease.
+When studying treatments for mild-to-moderate disease, the target population is 50% mild and 50% moderate.
+As shown in Figure 4, the ECE population for comparing treatments 1 and 2 includes all individuals with mild or moderate disease, since all can potentially be assigned to these two treatments, which exactly represents the population of interest – 50% mild and 50% moderate.
+In contrast, if only considering those participants assigned to treatments 1 or 2, or those within sub-study 1, the analysis population may arbitrarily depend on the randomization ratio between sub-studies.
+For instance, if the allocation ratio to sub-study 1 for the moderate disease population decreases from 1/2 to 1/4, the resulting population restricted to sub-study 1 shifts to 80% mild and 20% moderate.
+The SIMPLIFY trial has two sub-studies: one compares treatments 1 and 2 in a population on HS only or both HS and DA; the other compares treatments 1 and 3 in a population on DA only or both HS and DA.
+Participants on both HS and DA are first randomized between the two sub-studies using a 1:1 ratio, which was later adjusted to 3:1.
+As shown in Figure 1, the ECE population for comparing treatments 1 and 2 (i.e., assessing the effect of discontinuing HS versus continuing) includes all individuals on HS only or on both HS and DA, irrespective of the sub-study randomization ratio.
+This appropriately reflects the eligibility criteria for the treatments being compared.
+In contrast, a naive analysis restricted to participants assigned to the HS study targets an analysis population that differentially weights the HS-only and HS+DA groups, where the weights are determined by the randomization ratio among individuals on both HS and DA, as well as the timing of any changes to that ratio; further discussions are in section 6.
+Alternatively, an analysis that includes all participants assigned to treatments 1 or 2 may include DA-only individuals, making the resulting analysis population less relevant for evaluating the effect of discontinuing HS.
+In summary, the ECE population preserves the integrity of randomized comparison, is invariant to the randomization ratio or trial type, and aligns with the intention-to-treat principle (ICH E9,, 1998).
+It is an appropriate population for both traditional and platform trials, and is especially useful for providing a clear and logical framework for defining populations when evaluating treatment effects in platform trials.
+In contrast, any population defined by actual treatment assignments or sub-study allocations may not be reasonable, as they can be quite arbitrary and influenced by trial design or specific randomization probabilities.
+Additionally, individuals outside the ECE population have zero probability of being assigned to either treatment $j$ or $k$, due to ineligibility or unavailability of treatments, thereby necessitating extrapolation.
+In some situations, other populations, such as subgroups defined by specific characteristics, enrollment windows, or re-weighted populations from different enrollment windows, may be of interest.
+For example, in COVID-19 trials with rapidly evolving variants, the population enrolled during later windows, when newer variants are prevalent, may hold the greatest clinical relevance.
+The ECE population weights each enrollment window according to its natural size, but alternative weights that reflect clinical priorities can also be considered.
+The concept of ECE population can also be extended to comparisons involving more than two arms.
+For instance, the ECE population for comparing three treatments $j$, $k$, and $l$ is a population of all individuals for whom $\pi_{j}(Z)>0$, $\pi_{k}(Z)>0$, and $\pi_{l}(Z)>0$.
+In any case, we recommend starting with the ECE population as the baseline, then considering its subgroups or applying weighting to reflect the target population of interest.
+This approach offers clarity and specificity in defining the population of interest, ensuring that the population reflects the clinical question rather than having arbitrary dependence on design choices like randomization ratio or trial types.
+Moreover, the ECE population is crucial in other areas of statistical research, such as when using nonconcurrent controls to improve trial efficiency, as it provides a clear reference point for evaluating both efficiency gains and potential bias.
+These topics will be explored in future research.
+3.3 Treatment effect for the ECE population
+With the population now clearly defined, we propose the following estimand to define the treatment effect for two given treatments $j$ and $k$ within the ECE population:
 ${\bm{\vartheta}}_{jk}=\left(\begin{array}[]{c}\theta_{jk}\\ \theta_{kj}\end{array}\right)=\left(\begin{array}[]{c}E(Y^{(j)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\\ E(Y^{(k)}\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0)\end{array}\right),$ (1)
-where both subscripts $j$ and $k$ are used in $\bm{\vartheta}_{jk}$, $\theta_{jk}$, and $\theta_{kj}$ because both $\pi_{j}({Z})$ and $\pi_{k}({Z})$ are involved.
-In essence, a contrast of the estimand in (1) is an unconditional treatment effect for all concurrently eligible patients, averaging across all eligible patients and over time during which the enrollment window is open.
-Time can represent many background factors that may be changing although this is not explicitly recognized in estimand specification for traditional clinical trials.
-Our defined population of concurrently eligible patients is the largest population that can preserve the integrity of randomization when inferring treatment effects.
-Individuals outside this group have zero probability of being assigned to treatments $j$ or $k$, or both, due to their ineligibility or the unavailability of treatments, thereby necessitating extrapolation.
-This is the reason why we do not consider the population means $E(Y^{(j)})$ and $E(Y^{(k)})$ that are defined for the entire platform trial population.
-Conversely, our estimand in (1) represents the treatment effect for a broader population compared to just those enrolled in a specific sub-study.
-Our estimand in (1) truly respects the eligibility criteria for the given pair of treatments for comparison, rather than on the subset assigned to a particular sub-study.
-Throughout the paper, we use the notation convention that $a_{kj}$ is defined by swapping subscripts $j$ and $k$ in $a_{jk}$, e.g., $\theta_{jk}$ and $\theta_{kj}$ defined in (1).
-Based on the estimand ${\bm{\vartheta}}_{jk}$ in (1), the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$.
-Our approach to defining treatment effects is adaptable to situations where population characteristics other than means are considered, or when the treatment effect for a targeted subgroup is of primary interest.
+where $\theta_{jk}$ and $\theta_{kj}$ depend on $j$ and $k$ because the ECE population depends on them.
+A contrast of the estimand in (1) is an unconditional treatment effect for the ECE population, averaging across all eligible participants and over time during which the enrollment window is open.
+For example, the treatment effect of $j$ versus $k$ can be defined as the linear contrast $\theta_{jk}-\theta_{kj}$, the ratio $\theta_{jk}/\theta_{kj}$, or some other function of ${\bm{\vartheta}}_{jk}$.
+In the context of the SIMPLIFY trial, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS versus continuing in a population of all individuals taking only HS or both HS and DA.
+A similar interpretation applies to $\theta_{31}-\theta_{13}$.
+This estimand framework is readily extendable to settings with multiple treatment arms.
+Due to differences in the populations, $\theta_{jk}$ and $\theta_{kj}$ in (1) differ from $E(Y^{(j)})$ and $E(Y^{(k)})$, respectively, which are the means for the entire master protocol trial population and may include individuals who are non-concurrent or ineligible for treatments $j$ and $k$.
+Inference on $E(Y^{(j)})$ and $E(Y^{(k)})$ requires extrapolation.
+When these quantities are clinically of interest, one can apply methods to generalize the estimand in (1) to the entire trial population under additional assumptions (Degtiar and Rose,, 2023).
+Additionally, the estimand in (1) differs from the treatment effect for those assigned to specific arms (e.g., $E\{Y^{(j)}-Y^{(k)}\mid A\in(j,k)\}$), which, as shown by the formula, can depend on the distribution of $A$ and have problems such as arbitrary dependence on the randomization ratio.
+Similarly, the estimand in (1) also differs from the treatment effect for those assigned to specific sub-studies.
 4 Estimation Methods
+4.1 Analysis set
+The analysis set, distinct from the population used to define the estimand, refers to the sample used in statistical analysis.
+While the ECE population is uniquely specified, the analysis set is not unique and can include multiple options.
+To maximize efficiency, the analysis set can include all concurrently eligible individuals in the index set ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$.
+However, practical considerations may lead to using a subset of ${{{\cal I}}}_{jk}$.
+For example, when applying a partial blinding strategy (FDA, 2023b, ) or if other sub-studies remain blinded at the time of analysis, it may be preferable to restrict the analysis set to individuals within a specific sub-study.
+In these cases, the analysis set is a biased sample of the ECE population and requires appropriate corrections.
+A general discussion of analysis sets and the conditions they must satisfy is provided in Section S1 of the Supplement.
+To focus our discussions, in the following, we consider either the analysis set defined by ${{{\cal I}}}_{jk}$ (with covariate adjustment) or the subset ${{{\cal I}}}_{jk}\cap\{i:A_{i}=j\text{ or }A_{i}=k\}$ (without covariate adjustment), where $\cap$ denotes the intersection of two sets.
 4.2 Naive method
-After we define the estimand ${\bm{\vartheta}}_{jk}$ in (1), the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible patients defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z}),\pi_{k}({Z})$ may still vary over time and patient characteristics, leading to confounding.
-A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from patients with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$.
-While $\bar{Y}_{jk}$ is valid in traditional trials, it is typically biased under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome.
-Specifically, as shown in the Supplementary Material, as $n\to\infty$,
+After we define the estimand, the next challenge is estimating each component of ${\bm{\vartheta}}_{jk}$ in (1), because even among the concurrently eligible individuals defined by $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$, these non-zero probabilities $\pi_{j}({Z})$ and $\pi_{k}({Z})$ may still vary over time and individual characteristics, leading to confounding.
+A naive method is to estimate the component $\theta_{jk}$ by the sample mean $\bar{Y}_{jk}$ of $Y_{i}$’s from individuals with $\pi_{j}({Z}_{i})>0$, $\pi_{k}({Z}_{i})>0$, and $A_{i}=j$.
+However, $\bar{Y}_{jk}$ is typically biased for $\theta_{jk}$ under platform trials as it overlooks confounding by ${Z}$, i.e., the fact that $Z$ affects both treatment and outcome.
+Specifically, as shown in the Supplement, as $n\to\infty$,
 $\bar{Y}_{jk}\to\frac{E\{\pi_{j}({Z})E(Y^{(j)}\mid{Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}{E\{\pi_{j}({Z})\mid\pi_{j}({Z})>0,\pi_{k}({Z})>0\}}\quad\mbox{in probability}.$ (2)
 The limit of $\bar{Y}_{jk}$ in (2) is not equal to $\theta_{jk}$ in (1) unless either $\pi_{j}({Z})$ does not vary with ${Z}$ or $Y^{(j)}\perp{Z}$.
-Take the trial described in Section 2.1 and Figure 1 as an example, where $\pi_{j}({Z})$ varies with ${Z}$, the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited patients being healthier than those recruited earlier.
+Take the trial described in Figure 2 as an example, where $\pi_{j}({Z})$ varies with ${Z}$ (which does not occur in traditional clinical trials), the condition $Y^{(j)}\perp{Z}$ unrealistically rules out the influence of disease subtype on the outcome, as well as potential temporal effects, such as later recruited individuals being healthier than those recruited earlier.
 The problem with the naive estimator $\bar{Y}_{jk}$ arises from confounding by $\pi_{j}({Z})$, not from conditioning on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
-In traditional clinical trials, $\bar{Y}_{jk}$ is valid because $\pi_{j}({Z})$ does not vary with ${Z}$.
 4.3 Inverse probability weighting
-As noted in Section 4.1, the naive method is biased because it fails to account for confounding by $Z$.
-Under Assumption 1, since the true probability of receiving treatment $\pi_{j}(Z)$ is known, this confounding can be addressed using inverse probability weighting (IPW).
-The identification formula is (the proof is in the Supplementary Material):
-$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},\vspace{-3mm}$ (3)
+As noted in Section 4.2, the naive approach is biased because it does not account for confounding by $Z$.
+Under Assumption 1, since the true probability of receiving treatment, $\pi_{j}(Z)$, is known, this confounding can be addressed through inverse probability weighting (IPW).
+Notably, weighting by $\pi_{j}(Z)$ also helps correct for sample bias, as IPW uses only those assigned to treatment $j$, which might otherwise be a biased sample from the ECE population in platform trials (see the discussion at the end of Section 3.2).
+The IPW identification formula is (the proof is in the Supplement):
+$\theta_{jk}=E\bigg{\{}\frac{I(A=j)Y}{\pi_{j}({Z})}\ \Big{|}\ \pi_{j}({Z})>0,\pi_{k}({Z})>0\bigg{\}},$ (3)
 where $I(\cdot)$ denotes the indicator function.
 Formula (3) naturally suggests the following IPW estimator of $\theta_{jk}$, which is widely used in survey sampling (Horvitz and Thompson,, 1952) and causal inference (Rosenbaum and Rubin,, 1983),
-$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},\vspace{-3mm}$
+$\hat{\theta}_{jk}^{\rm\,ipw}=\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})},$
 where ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ and $n_{jk}$ is the number of elements in ${{{\cal I}}}_{jk}$.
-This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971; Robins et al.,, 2000) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
+This IPW estimator may be improved by the following stabilized IPW (SIPW) (Hájek,, 1971) estimator, which normalizes the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$ to sum to 1:
 $\hat{\theta}_{jk}^{\rm\,sipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)Y_{i}}{\pi_{j}({Z}_{i})}.$
@@
 The IPW and SIPW methods utilize some covariate information through weighting with $\pi_{j}({Z})$.
-To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2023), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect.
-This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (ICH E9,, 1998; EMA,, 2015; FDA,, 2021).
-Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may contain or overlap with ${Z}$.
-We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a potentially misspecified working model.
-For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y_{i}$ on $\bm{X}$ using data where $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$.
+To improve efficiency, we can adjust for ${Z}$ and other baseline covariates using a model-assisted approach (Tsiatis et al.,, 2008; Ye et al.,, 2023; Bannick et al.,, 2025), which incorporates these covariates through a working model between the outcome and covariates and produces an asymptotically unbiased estimator even if the working model is incorrect.
+This robustness makes model-assisted approaches increasingly popular and recommended by the regulatory agencies (FDA, 2023a, ).
+Specifically, let $\bm{X}\subset\bm{W}$ denote the observed vector of baseline covariates for adjustment, which may overlap with ${Z}$.
+We estimate $E(Y^{(j)}\mid\bm{X},\pi_{j}({Z})>0,\pi_{k}({Z})>0,A=j)$ by $\hat{\mu}_{jk}(\bm{X})$, using a working model that may be misspecified.
+For continuous outcomes, $\hat{\mu}_{jk}(\bm{X})$ is usually from least squares regression of $Y^{(j)}$ on $\bm{X}$ using data with $\pi_{j}({Z})>0,\pi_{k}({Z})>0$ and $A=j$.
 For binary outcomes, $\hat{\mu}_{jk}(\bm{X})$ is often obtained from a logistic working model.
-After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for covariate $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
+After $\hat{\mu}_{jk}(\cdot)$ is obtained, we can adjust for $\bm{X}$ using the augmented inverse probability weighting (AIPW) (Robins et al.,, 1994), resulting in the following AIPW estimator of $\theta_{jk}$,
 $\hat{\theta}_{jk}^{\rm\,aipw}=$ $\ \frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).\vspace{-3mm}$
-Interestingly, in its last term, the AIPW estimator leverages covariate information from all concurrently eligible patients in ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$, including those not receiving treatments $j$ or $k$.
-For example, in Figure 1, when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2 using AIPW, the estimator also incorporates covariate data from subtype A patients assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2.
-These covariate data are not used in the IPW and SIPW estimators.
-Thus, AIPW offers a robust method to pool data from all concurrently eligible patients, thereby further extracting the potential of efficiency gain through platform trials.
+Interestingly, in its last term, the AIPW estimator leverages covariate information from all individuals within ${{{\cal I}}}_{jk}$, including those not receiving treatments $j$ or $k$.
+For example, in Figure 2(a), when estimating the effect of treatment 3 versus treatment 1 at the end of enrollment window 2, the AIPW also incorporates covariate data from subtype A individuals assigned to treatment 2 (in sub-study 1) and treatment 4 (in sub-study 3) during enrollment windows 1 and 2.
+These covariate data are not used in the IPW or SIPW estimators.
+Thus, AIPW offers a robust method to pool data from all individuals within ${{{\cal I}}}_{jk}$, thereby further extracting the potential of efficiency gain through platform trials.
 Similarly, one can use the stabilized AIPW (SAIPW) estimator by normalizing the weights $I(A_{i}=j)/\pi_{j}({Z}_{i})$ for $i\in{{{\cal I}}}_{jk}$, resulting in the following SAIPW estimator of $\theta_{jk}$,
 $\hat{\theta}_{jk}^{\rm\,saipw}=\bigg{\{}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)}{\pi_{j}({Z}_{i})}\bigg{\}}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}}{\pi_{j}({Z}_{i})}+\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
-Both AIPW and SAIPW estimators are asymptotically normal (see Theorem 1 in Section 5).
-As discussed in Section 5, whether to stabilize the weights in the AIPW approach often results in asymptotically equivalent estimators.
+Both AIPW and SAIPW estimators are asymptotically normal (Theorem 1 in Section 5).
+Note that when $Z_{i}$ is discrete, $\pi_{j}(Z_{i})$ can be estimated using sample proportions and used in place of the true value $\pi_{j}(Z)$ in any of the above weighting estimators.
 4.4 Post-stratification
-An alternative method to account for the varying $\pi_{j}({Z})$ is to stratify subjects based on the values that $\pi_{j}({Z})$ can take.
-This approach is known as post-stratification in sample surveys (Fuller,, 2009) or sub-classification in the causal inference literature (Rosenbaum and Rubin,, 1984).
-Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible patients ${{{\cal I}}}_{jk}=\{i:\pi_{j}({Z}_{i})>0,\pi_{k}({Z}_{i})>0\}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$.
+When $\pi_{j}({Z})$ takes discrete values, an alternative method to account for the varying $\pi_{j}({Z})$ is to stratify individuals based on the values that $\pi_{j}({Z})$ can take.
+This approach is known as post-stratification in sample surveys (Fuller,, 2009).
+Specifically, for a given pair $j$ and $k$, we divide all concurrently eligible individuals in ${{{\cal I}}}_{jk}$ into a finite number of strata ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$, where $H_{jk}$ is the number of strata, such that within each stratum ${{\cal I}}_{jk}^{(h)}$, the values of $\pi_{j}({Z}_{i})$ and $\pi_{k}({Z}_{i})$ are constant, denoted as $\pi_{j}({\cal I}_{jk}^{(h)})$ and $\pi_{k}({\cal I}_{jk}^{(h)})$, respectively, for all $i$ in ${{\cal I}}_{jk}^{(h)}$.
 The resulting post-stratification (PS) estimator of $\theta_{jk}$ is
 $\hat{\theta}_{jk}^{\rm\,ps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)Y_{i}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})\bar{Y}_{j}({\cal I}_{jk}^{(h)}),$
-where $n({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of patients in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$.
-This estimator does not need stabilizing, as the sum of weights is
-$\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{{\cal I}}}_{jk}^{(h)}}I(A_{i}=j)=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}n({{\cal I}}_{jk}^{(h)})=1.$
-Compared with the SIPW, the PS requires an extra step of constructing strata for every pair $j$ and $k$.
+where $n({{\cal I}}_{jk}^{(h)})$ is the number of individuals in ${{\cal I}}_{jk}^{(h)}$, $n_{j}({{\cal I}}_{jk}^{(h)})$ is the number of individuals in ${{\cal I}}_{jk}^{(h)}$ with $A_{i}=j$, and $\bar{Y}_{j}({\cal I}_{jk}^{(h)})$ is the sample mean of $Y_{i}$’s for units in ${\cal I}_{jk}^{(h)}$ with $A_{i}=j$.
+This estimator does not need stabilizing, as the sum of weights in $\hat{\theta}_{jk}^{\rm\,ps}$ is equal to 1.
 The idea of PS also appeared in Marschner and Schou, (2022), but their approach of stratifying by enrollment window is inadequate unless ${Z}$ consists solely of enrollment time.
-For instance, in Figure 1, ${Z}$ includes both the disease subtype and the enrollment window; in the SIMPLIFY example in Figure 2, $Z$ is the indicator of HS and/or DA and there is only a single enrollment window.
 On the other hand, stratifying according to all joint levels of enrollment window and disease subtype is unnecessary.
 Our proposed method of stratification ensures that $\pi_{j}({Z})$ and $\pi_{k}({Z})$ remain constant within each post-stratum, which is exactly what is required to handle the issue of varying $\pi_{j}({Z})$ and $\pi_{k}({Z})$.
-It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section 7.1.
+It is important to note that post-stratification with redundant strata may lead to excessively small strata, adversely affecting the finite sample performance of PS estimator or its variance estimator; this is further explored through simulation results and discussed in Section S2 of the Supplement.
 The PS approach utilizes covariate information by stratifying through ${{\cal I}}_{jk}^{(1)},...,{{\cal I}}_{jk}^{(H_{jk})}$.
 To adjust for additional covariates, we apply AIPW separately within each stratum and then combine the results using weighted averages.
-This gives the following adjusted post-stratification (APS) estimator of $\theta_{jk}$,
+This gives the adjusted post-stratification (APS) estimator of $\theta_{jk}$,
 $\hat{\theta}_{jk}^{\rm\,aps}=\frac{1}{n_{jk}}\sum_{h=1}^{H_{jk}}\frac{n({{\cal I}}_{jk}^{(h)})}{n_{j}({{\cal I}}_{jk}^{(h)})}\sum_{i\in{{\cal I}}_{jk}^{(h)}}I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}+\frac{1}{n_{jk}}\sum_{i\in{{\cal I}}_{jk}}\hat{\mu}_{jk}(\bm{X}_{i}).$
-Asymptotically, the PS and APS estimators are shown to be equivalent to certain types of SAIPW estimators, as detailed in Section 5.
-4.4 Summary of estimators
-[…]
 5 Asymptotic Theory
@@
 For any given $j$ and $k$, there exists a function $\mu_{jk}(\cdot)$ with finite $E_{X}\{\mu_{jk}^{2}(\bm{X})\}$ such that, as $n\to\infty$, $E_{X}\{\hat{\mu}_{jk}(\bm{X})-\mu_{jk}(\bm{X})\}^{2}\to 0$ in probability with respect to the randomness of $\hat{\mu}_{jk}(\cdot)$ as a function of data, where $E_{X}$ is the expectation with respect to $\bm{X}$.
-If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in S1 of the Supplementary Material.
+If $\hat{\mu}_{jk}(\cdot)$ is not from a finite-dimensional parametric model, then $\mu_{jk}(\cdot)$ and $\hat{\mu}_{jk}(\cdot)$ also need to satisfy the Donsker condition stated in Section S1 of the Supplement.
 Theorem 1 shows that all six estimators described in Section 4 are consistent in estimating ${\bm{\vartheta}}_{jk}$ and are asymptotically normal under minimal statistical assumptions.
 In what follows, $E_{jk}=E_{kj}$, ${\rm Var}_{jk}={\rm Var}_{kj}$, and ${\rm Cov}_{jk}={\rm Cov}_{kj}$ denote respectively the expectation, variance, and covariance conditioned on $\pi_{j}({Z})>0$ and $\pi_{k}({Z})>0$.
+Similarly, $E_{jk}(\cdot\mid S)=E_{kj}(\cdot\mid S)$, ${\rm Var}_{jk}(\cdot\mid S)={\rm Var}_{kj}(\cdot\mid S)$, and ${\rm Cov}_{jk}(\cdot\mid S)={\rm Cov}_{kj}(\cdot\mid S)$ denote these quantities further conditioned on $S$, where $S$ is the covariate for post-stratification, i.e., for individual $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$.
 Theorem 1.
-Under Assumption 1, and Assumption 2 if adjusting for covariates, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$ whose explicit expression is given in each of the following specific case.
+Under Assumption 1, and Assumption 2 if covariates are adjusted, for fixed $j$ and $k$, and $\star\in\{{\rm ipw,sipw,aipw,saipw,ps,aps}\}$, $n_{jk}^{1/2}({\hat{\bm{\vartheta}}^{\star}_{jk}}-{\bm{\vartheta}}_{jk})$ converges in distribution as $n\to\infty$ to the bivariate normal with mean vector 0 and covariance matrix $\bm{\Sigma}^{\star}_{jk}$, where $\hat{\bm{\vartheta}}_{jk}^{\star}=(\hat{\theta}_{jk}^{\star},\hat{\theta}_{kj}^{\star})^{T}$, $\bm{a}^{T}$ is the vector transpose of $\bm{a}$, and the explicit expression of $\bm{\Sigma}^{\star}_{jk}$ is given in each of the following specific case.
 (a) For the
@@
 $\bm{\Sigma}^{\rm ps}_{jk}=E_{jk}\Big{[}{\rm diag}\left\{{\textstyle\frac{{\rm Var}_{jk}(Y^{(j)}\mid S)}{\pi_{j}(S)},\,\frac{{\rm Var}_{jk}(Y^{(k)}\mid S)}{\pi_{k}(S)}}\right\}\Big{]}+{\rm Var}_{jk}\{E_{jk}(Y^{(j)}\mid S),\,E_{jk}(Y^{(k)}\mid S)\},\vspace{-2mm}$
-where $S$ is the covariate for post-stratification, i.e., for patient $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplementary Material.
+where $S$ is the covariate for post-stratification, i.e., for individual $i$, $S_{i}={\cal I}_{jk}^{(h)}$ if and only if $i\in{\cal I}_{jk}^{(h)}$, and $\pi_{j}(S)=P(A=j\mid S)=\pi_{j}({Z})$ by Lemma S1 in the Supplement.
 (f) For the
@@
 where $\bm{\Lambda}_{jk}(S)$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ replaced by $\lambda_{jk}(S)=2{\rm Cov}_{jk}\{Y^{(j)},\mu_{jk}(\bm{X})\mid S\}-{\rm Var}_{jk}\{\mu_{jk}(\bm{X})\mid S\}$ and $c_{jk}(S)={\rm Cov}_{jk}\{Y^{(j)},$ $\mu_{kj}(\bm{X})\mid S\}$ $+\,{\rm Cov}_{jk}\{Y^{(k)},\mu_{jk}(\bm{X})\mid S\}-{\rm Cov}_{jk}\{\mu_{jk}(\bm{X}),\mu_{kj}(\bm{X})\mid S\}$, respectively.
-All technical proofs are given in the Supplementary Material.
-The results in Theorem 1 for estimators with covariate adjustment hold regardless of whether the working model is correct or not.
-The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators in Section 6.
+All technical proofs are given in the Supplement.
+The results in Theorem 1 and the efficiency comparison in Section 5.2 are new to the literature and differ from existing results for IPW or AIPW estimators using alternative weights, such as estimated propensity scores (Robins et al.,, 1994; Lunceford and Davidian,, 2004).
+We also show that the AIPW and SAIPW estimators with $\mu_{jk}=E_{jk}(Y^{(j)}\mid\bm{X})$ and $\mu_{kj}=E_{jk}(Y^{(k)}\mid\bm{X})$ are asymptotically equivalent and both attain semiparametric efficiency.
+The proof is provided in Section S3 of the Supplement.
+Furthermore, our results include PS and APS estimators and hold regardless of whether the working model is correctly specified.
+The asymptotic covariance matrices are explicitly given, which motivates robust variance estimators presented in Section S1 of the Supplement.
 5.2 Efficiency comparison
-As the six estimators proposed in Section 4 have different asymptotic covariance matrices, we compare the asymptotic relative efficiency between pairs of these estimators in the subsequent corollaries.
+We compare the asymptotic relative efficiency between pairs of six estimators in the subsequent corollaries.
 The first two results focus on comparing the efficiency of stabilized versus unstabilized estimators.
 Corollary 1 (Comparison of IPW and SIPW).
-Under Assumption 1, the difference in asymptotic covariance matrices of the IPW and SIPW estimators is equal to
-$\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$
-$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.$
-Since the matrix $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ in Corolary 3 is positive semidefinite (shown in S6 of the Supplementary Material) but ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ may have arbitrary signs, the asymptotic relative efficiency of IPW versus SIPW is indefinite.
-When ${\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}}$ and ${\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}$ are both non-negative, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}$ is positive semidefinite and, hence, SIPW always has equal or better asymptotic efficiency than IPW in estimating any differentiable function of $\bm{\vartheta}_{jk}$.
-On the other hand, an example of estimating $\theta_{jk}-\theta_{kj}$ is shown in S6 of the Supplement Material in which IPW is asymptotically more efficient than SIPW.
-Nonetheless, SIPW is often more efficient in simulations and real data applications (see Section 7).
-SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change the SIPW estimator $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect the IPW analogue $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023).
+Under Assumption 1, $\bm{\Sigma}_{jk}^{\rm\,ipw}-\bm{\Sigma}_{jk}^{\rm\,sipw}=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)},\frac{\theta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)},\frac{\theta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}$.
+Although $E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\theta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\theta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\vartheta}_{jk}\bm{\vartheta}_{jk}^{T}$ is positive semidefinite, the asymptotic relative efficiency of IPW versus SIPW is indefinite because the covariances may have arbitrary signs; see Section S3 of the Supplement for more discussions.
+Nonetheless, SIPW is often more efficient in simulations (see Section S2 of the Supplement).
+SIPW is also preferable due to its invariance to outcome shifts for estimating $\theta_{jk}-\theta_{kj}$, i.e., adding a constant to every outcome does not change $\hat{\theta}_{jk}^{\rm sipw}-\hat{\theta}_{kj}^{\rm sipw}$ but does affect $\hat{\theta}_{jk}^{\rm ipw}-\hat{\theta}_{kj}^{\rm ipw}$ (Ding,, 2023).
 For these reasons, we do not consider IPW in the remaining efficiency comparisons.
 Corollary 2 (Comparison of AIPW and SAIPW).
-Under Assumptions 1-2, the difference in asymptotic covariance matrices of the AIPW and SAIPW estimators is equal to
-$\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}$ $=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}$
-$\quad+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},\,{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}.\vspace{-2mm}$
-Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition
-$\delta_{jk}=\delta_{kj}=0,\vspace{-3mm}\vspace{-2mm}$ (4)
-which is known as the prediction unbiasedness condition (Guo and Basse,, 2021) at the population level, where $\delta_{jk}=E_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X})\}$ and $\delta_{kj}=E_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X})\}$.
-Condition (4) holds when the working models used for covariate adjustment are correctly specified.
-It also holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified.
-Later discussions about adjustment (8) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet condition (4), it can be readily modified by (8) to satisfy this requirement.
+Under Assumptions 1-2, $\bm{\Sigma}_{jk}^{\rm\,aipw}-\bm{\Sigma}_{jk}^{\rm\,saipw}=E_{jk}\Big{[}{\rm diag}\Big{\{}{\textstyle\frac{\delta_{jk}^{2}}{\pi_{j}({Z})},\,\frac{\delta_{kj}^{2}}{\pi_{k}({Z})}}\Big{\}}\Big{]}-\bm{\delta}_{jk}\bm{\delta}_{jk}^{T}+2\,{\rm diag}\Big{\{}{\textstyle{\rm Cov}_{jk}\big{\{}Y^{(j)}-\mu_{jk}(\bm{X}),\frac{\delta_{jk}}{\pi_{j}({Z})}\big{\}},}$ ${\textstyle{\rm Cov}_{jk}\big{\{}Y^{(k)}-\mu_{kj}(\bm{X}),\frac{\delta_{kj}}{\pi_{k}({Z})}\big{\}}}\Big{\}}$.
+Corollary 2 shows that the AIPW and SAIPW estimators are asymptotically equivalent, i.e., $\bm{\Sigma}_{jk}^{\rm\,aipw}=\bm{\Sigma}_{jk}^{\rm\,saipw}$, under the condition $\delta_{jk}=\delta_{kj}=0$, which holds for many widely-used regression models, such as generalized linear models with canonical links, even when the working models are not correctly specified.
+Later discussions about adjustment (7) shows that, even if the initial $\hat{\mu}_{jk}(\bm{X})$ does not meet this condition, it can be readily modified by (7) to satisfy this requirement.
 Therefore, AIPW and SAIPW are usually asymptotically equivalent and we do not include AIPW in the rest of the efficiency comparisons.
-The next two results show when covariate adjustment leads to an efficiency gain. Corollary 3 compares the two stabilized IPW estimators, with or without covariate adjustment.
+The next two results show when covariate adjustment leads to an efficiency gain.
 Corollary 3 (Comparison of SIPW and SAIPW).
@@
 ${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{jk}(\bm{X})\mid S\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{kj}(\bm{X})\mid S\}=0,$ (5)
-${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{kj}(\bm{X})\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{jk}(\bm{X})\}=0,$ (7)
+${\rm Cov}_{jk}\{Y^{(j)}-\mu_{jk}(\bm{X}),\,\mu_{kj}(\bm{X})\}={\rm Cov}_{jk}\{Y^{(k)}-\mu_{kj}(\bm{X}),\,\mu_{jk}(\bm{X})\}=0,\vspace{-2mm}$ (6)
 where $S$ is the covariate for post-stratification as defined in Theorem 1(e), the SAIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ is asymptotically more efficient than the SIPW estimator $\hat{\bm{\vartheta}}_{jk}^{\rm\,sipw}$, i.e.,
@@
 When covariate adjustment for $\bm{X}$ is applied, the SAIPW estimator is not necessarily more efficient than the SIPW estimator if the working models are incorrect.
-Conditions (5)-(7) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the platform trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent.
-Actually, condition (5) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
+Conditions (4)-(6) in Corollary 3 are sufficient for guaranteeing an efficiency gain after covariate adjustment in estimating any differentiable function of $\bm{\vartheta}$, except in the trivial scenario where one of $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ is constant or $\pi_{j}(S)+\pi_{k}(S)=1$ (there are only two treatments in the master protocol trial) and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent.
+Condition (4) holds without loss of generality since it can always be achieved by replacing $\hat{\mu}_{jk}(\bm{X})$ with
 $\tilde{\mu}_{jk}(\bm{X})=\hat{\mu}_{jk}(\bm{X})+\frac{1}{n_{jk}(S)}\sum_{i\in{\cal I}_{jk},A_{i}=j,S_{i}=S}\{Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i})\},$ (7)
@@
 Otherwise, they require certain construction of $\hat{\mu}_{jk}(\bm{X})$ and $\hat{\mu}_{kj}(\bm{X})$ that has (5)-(6) embedded.
-For example, (5)-(7) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2023); see S10 in the Supplement Material for details.
+For example, (4)-(6) hold if applying a linear ANHECOVA working model that includes $S$ and its interaction with $\bm{X}$ (Ye et al.,, 2023), or applying joint calibration when a non-linear working model is used (Bannick et al.,, 2025); see S3 in the Supplement for details.
 In fact, (4)-(6) always hold for the SAIPW estimator with $\bm{X}=S$, i.e., the SAIPW estimator with $\hat{\mu}_{jk}(\bm{X})=\bar{Y}_{j}(S)$, which is the sample mean of $Y_{i}$’s for $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}=S$.
@@
 Therefore, the SAIPW estimator with $\bm{X}=S$ is asymptotically more efficient than the SIPW estimator according to Corollary 3 and it is used in later corollaries to bridge the comparison between IPW-type and PS-type estimators.
-Corollary 4 compares the two post-stratification estimators, with or without additional adjustment for covariates.
+Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in some trivial scenarios.
 Corollary 4 (Comparison of PS and APS).
@@
 is positive definite, unless either one of ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ and ${\rm Var}_{jk}\{\mu_{kj}(\bm{X})\mid S\}$ is 0 or $\pi_{j}(S)+\pi_{k}(S)=1$ and the correlation between $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ conditioned on $S$ is ±$\pm 1$, in which case $\bm{\Sigma}_{jk}^{\rm\,ps}-\bm{\Sigma}_{jk}^{\rm\,aps}$ is positive semidefinite.
-Corollary 4 indicates a guaranteed efficiency gain after covariate adjustment on top of post-stratification, except in the trivial scenario where either no covariate adjustment beyond $S$ in one of the treatment arms or $\pi_{j}(S)+\pi_{k}(S)=1$ and $\mu_{jk}(\bm{X})$ and $\mu_{kj}(\bm{X})$ are linearly dependent conditioned on $S$.
 To compare the PS estimator with the IPW-type estimator, intuitively the PS estimator is asymptotically more efficient than the SIPW estimator because the former utilizes $S$ as a covariate in post-stratification whereas the latter only partially uses covariates through weighting.
 For a fair comparison, the PS estimator should be compared with the SAIPW estimator with $\bm{X}=S$, i.e., both estimators fully adjust for the discrete covariate $S$.
-This leads to the following result.
+The following result shows the asymptotic equivalence between these two estimators.
 Corollary 5 (Comparison of PS and SAIPW with $\bm{X}=S$).
-Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,ps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$ when $\bm{X}=S$.
-Corollary 5 states that the PS estimator that stratifies on $S$ is asymptotically equivalent to the SAIPW estimator that adjusts for $S$ (as well as the corresponding AIPW estimator due to its equivalence with the SAIPW when (4) holds).
-The next corollary compares APS and SAIPW with the same covariate adjustment.
+Under Assumption 1, $\hat{\bm{\vartheta}}_{jk}^{\rm\,ps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ with $\bm{X}=S$ have the same asymptotic covariance matrix.
+The last corollary compares APS and SAIPW with the same covariate adjustment.
 It demonstrates that when both estimators adjust for $\bm{X}$ using a working model that satisfies condition (4), the APS estimator is asymptotically equivalent to the SAIPW estimator.
 Corollary 6 (Comparison of APS and SAIPW).
-Under Assumptions 1-2 and condition (5), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix, i.e., $\bm{\Sigma}_{jk}^{\rm\,aps}=\bm{\Sigma}_{jk}^{\rm\,saipw}$.
-The last corollary completes the comparison around SAIPW with $\bm{X}=S$ and shows that under (5)-(7), adjusting for more covariates beyond just $S$ leads to additional asymptotic efficiency gain.
-Corollary 7 (Comparison of SAIPW and SAIPW with $\bm{X}=S$).
-Under Assumptions 1-2 and conditions (5)-(7), SAIPW with $S\subset\bm{X}$ is more efficient than SAIPW with $\bm{X}=S$.
+Under Assumptions 1-2 and condition (4), $\hat{\bm{\vartheta}}_{jk}^{\rm\,aps}$ and $\hat{\bm{\vartheta}}_{jk}^{\rm\,saipw}$ have the same asymptotic covariance matrix.
+It follows from Corollaries 4-6 that, under (4)-(6), SAIPW with $\bm{X}\supset S$ is more efficient than SAIPW with $\bm{X}=S$; see Section S3 in the Supplement for a proof.
 All the efficiency comparisons are summarized in Figure 5.
-Figure 3: Efficiency comparisons among all robust estimators.
-6 Robust Variance Estimation
-To robustly assess variability and conduct large sample inference, it is essential to obtain consistent estimators of the asymptotic covariance matrices in Theorem 1 for various estimators of $\bm{\vartheta}_{jk}$.
-These estimators should remain consistent whether the working models are correct or misspecified.
-For IPW-type estimators, we propose the following estimators of covariance matrices:
-$\hat{\bm{\Sigma}}_{jk}^{\rm\,ipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{\cal I}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)Y_{i}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)Y_{i}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}-\hat{\bm{\vartheta}}_{jk}\hat{\bm{\vartheta}}_{jk}^{T},$
-$\hat{\bm{\Sigma}}_{jk}^{\rm\,sipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)(Y_{i}-\hat{\theta}_{jk})^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)(Y_{i}-\hat{\theta}_{kj})^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\},$
-$\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk}-\hat{\bm{\delta}}_{jk}\hat{\bm{\delta}}_{jk}^{T},$
-$\hat{\bm{\Sigma}}_{jk}^{\rm\,saipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})-\hat{\delta}_{jk}\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})-\hat{\delta}_{kj}\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk},$
-where $\hat{\bm{\vartheta}}_{jk}=(\hat{\theta}_{jk},\hat{\theta}_{kj})^{T}$, $\hat{\theta}_{jk}$ is any estimator of $\theta_{jk}$, $\hat{\bm{\delta}}_{jk}=(\hat{\delta}_{jk},\hat{\delta}_{kj})^{T}$, $\hat{\delta}_{jk}=n_{jk}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\!I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}/\pi_{j}({Z}_{i})$, $\hat{\bm{\Lambda}}_{jk}$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ respectively estimated by $\hat{\lambda}_{jk}=2\hat{q}_{jk}^{(j)}+\hat{\sigma}^{2}_{jk}$ and $\hat{c}_{jk}=\hat{q}_{kj}^{(j)}+\hat{q}_{jk}^{(k)}+\hat{q}_{jk}$, $\hat{q}_{jk}^{(j)}$ is the sample covariance of $(Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{jk}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$ and $A_{i}=j$, $\hat{q}_{kj}^{(j)}$ is $\hat{q}_{jk}^{(j)}$ with $\hat{\mu}_{jk}(\bm{X}_{i})$ switched to $\hat{\mu}_{kj}(\bm{X}_{i})$, $\hat{q}_{jk}^{(k)}$ is $\hat{q}_{jk}^{(j)}$ with $Y_{i}^{(j)}$ switched to $Y_{i}^{(k)}$, $\hat{\sigma}^{2}_{jk}$ and $\hat{\sigma}^{2}_{kj}$ are the diagonal entries and $\hat{q}_{jk}$ is the off-diagonal entry of the sample covariance matrix of $(\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{kj}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$.
-For the PS and APS estimators, we propose the following estimators of $\bm{\Sigma}_{jk}^{\rm\,ps}$ and $\bm{\Sigma}_{jk}^{\rm\,aps}$,
-$\hat{\bm{\Sigma}}_{jk}^{\rm\,ps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,{\rm diag}\!\left\{{\textstyle\frac{\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\sigma}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Gamma}}_{jk}$
-$\hat{\bm{\Sigma}}_{jk}^{\rm\,aps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,\Big{[}{\rm diag}\!\left\{{\textstyle\frac{\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\tau}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})\Big{]}+\hat{\bm{\Gamma}}_{jk},$
-where $\hat{\pi}_{j}({\cal I}_{jk}^{(h)})=n_{j}({\cal I}_{jk}^{(h)})/n({\cal I}_{jk}^{(h)})$, $\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $Y_{i}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ estimated respectively by $\hat{\lambda}_{jk}({\cal I}_{jk}^{(h)})=2\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})+\hat{\sigma}_{jk}^{2}({\cal I}_{jk}^{(h)})$ and $\hat{c}_{jk}({\cal I}_{jk}^{(h)})=\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}^{(k)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}({\cal I}_{jk}^{(h)})$, $\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{\sigma}^{2}_{jk}({\cal I}_{jk}^{(h)})$, and $\hat{q}_{jk}({\cal I}_{jk}^{(h)})$ are respectively the same as $\hat{q}_{jk}^{(j)}$, $\hat{q}_{kj}^{(j)}$, $\hat{\sigma}^{2}_{jk}$, and $\hat{q}_{jk}$ previously defined in $\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}$ but with the index $i$ restricted to $i\in{\cal I}_{jk}^{(h)}$, $\hat{\bm{\Gamma}}_{jk}$ is the sample covariance matrix of $(\bar{Y}_{j}(S_{i}),\bar{Y}_{k}(S_{i}))$’s for $i\in{\cal I}_{jk}$, and $\bar{Y}_{j}(S_{i})$ is the sample mean of $Y_{i}$’s with $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}={\cal I}_{jk}^{(h)}$.
-7 Empirical Results
-7.1 Simulation study
-We conduct a simulation study based on the first three enrollment windows outlined in Figure 1 to examine the finite-sample performance of all estimators presented in Section 4.
-In the simulation, the observed baseline covariate vector is $\bm{W}=(X_{c},X_{b},Z_{\rm sub},Z_{\rm win})$, where $X_{c}$ is a continuous covariate uniformly distributed over the interval $(-3,3)$, $X_{b}$ is a binary covariate with $P(X_{b}=1)=0.5$, $Z_{\rm sub}$ is a binary indicator for disease subtype with $P(Z_{\rm sub}=1)=0.8$, and $Z_{\rm win}\in\{1,2,3\}$ denotes the enrollment window and satisfies
-$P(Z_{\rm win}=t\mid X_{c},X_{b},Z_{\rm sub},U)=\frac{\exp(Q_{t})}{\exp(Q_{1})+\exp(Q_{2})+\exp(Q_{3})},\quad t=1,2,3,$
-$Q_{1}=0.5+X_{c}+2X_{b}-Z_{\rm sub}+U$, $Q_{2}=1+2X_{c}+X_{b}-Z_{\rm sub}+U$, $Q_{3}=-0.5+X_{c}+X_{b}+Z_{\rm sub}+U$, and the unobserved baseline variable $U$ follows a standard normal distribution and is independent of $X_{c}$, $X_{b}$, and $Z_{\rm sub}$.
-The potential outcomes are generated as follows: $Y^{(1)}=1+X_{c}+X_{b}+Z_{\rm sub}+U+\epsilon_{1},$ $Y^{(2)}=1+X_{c}^{2}+X_{b}+Z_{\rm sub}+U+\epsilon_{2}$, $Y^{(3)}=3+X_{c}X_{b}+Z_{\rm sub}+U+\epsilon_{3}$, $Y^{(4)}=2+X_{c}Z_{\rm sub}-X_{b}+2U+\epsilon_{4}$, where $\epsilon_{j}$’s are mutually independent, each follows a standard normal distribution and is independent of $\bm{W}$ and $U$.
-Note that $Y^{(j)}$’s are correlated with $X_{c}$, $X_{b}$, $Z_{\rm sub}$, $U$, as well as indirectly with the enrollment time $Z_{\rm win}$.
-Table 1: Assignment probabilities to each sub-study and treatment
-${Z}$ Sub-study Treatment
-$Z_{\rm win}$ $Z_{\rm sub}$ 1 2 3 $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$ $\pi_{4}({Z})$
-1 1 0.4 0.6 0 0.50 0.20 0.30 0
-1 0 1 0 0 0.50 0.50 0 0
-2 1 0.3 0.3 0.4 0.50 0.15 0.15 0.20
-2 0 1 0 0 0.50 0.50 0 0
-3 1 0.4 0 0.6 0.50 0.20 0 0.30
-3 0 1 0 0 0.50 0.50 0 0
-Patients are first randomized into one of three sub-studies (labeled 1-3) depending on their enrollment windows and disease subtypes, and then to treatments within the assigned sub-study.
-Sub-study 1 has treatments 1 and 2, sub-study 2 has treatments 1 and 3, and sub-study 3 has treatments 1 and 4.
-Within each sub-study, treatments are randomized in a 1:1 ratio.
-The assignment probabilities are in Table 1.
-For the total sample size $n$ in the platform trial, we consider 500 and 1,000.
-When $n=500$, the expected numbers of patients per treatment are 123, 51, and 76 in sub-studies 1-3, respectively.
-We consider the following estimators of the linear contrast $\theta_{j1}-\theta_{1j}$, for $j=2,3,4.$
-  1. (i)
-    Four IPW-type estimators.
-We consider the IPW, SIPW, and SAIPW estimators described in Section 4.2.
-For SAIPW, $\hat{\mu}_{jk}$ is obtained from fitting a linear regression of $Y$ on $(X_{c},X_{b},Z_{\rm sub})$, using data from treatment arm $j$ in ${\cal I}_{jk}$.
-Since condition (4) holds in this simulation, the AIPW estimator is numerically almost the same as the SAIPW estimator and thus is omitted.
-We add the SAIPW estimator with $\bm{X}=S$, denoted by SAIPW($S$), to check its equivalence with the PS estimator as discussed in Corollary 5.
-Note that the working models are correctly specified only for treatment arm 1, as the true models for the other arms are not linear in $(X_{c},X_{b},Z_{\rm sub})$.
-  2. (ii)
-    Four PS-type estimators. We consider the following strata for PS and APS:
-    Estimating ${\small\begin{array}[]{ll}\mbox{Estimating $\theta_{21}-\theta_{12}$:}&{\cal I}_{12}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.5\,(Z_{{\rm sub},i}=0)\},\\ &{\cal I}_{12}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.2\,(Z_{{\rm win},i}=1\ \mbox{or}\ 3,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{12}^{(3)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{31}-\theta_{13}$:}&{\cal I}_{13}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.3\,(Z_{{\rm win},i}=1,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{13}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{41}-\theta_{14}$:}&{\cal I}_{14}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.2\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{14}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.3\,(Z_{{\rm win},i}=3,Z_{{\rm sub},i}=1)\}.\end{array}}\vspace{2mm}$
-    The same working models $\hat{\mu}_{jk}$ for SAIPW are used for APS.
-For estimating $\theta_{21}-\theta_{12}$, we additionally include the PS and APS estimators that stratify on all six joint levels of ${Z}=(Z_{\rm win},Z_{\rm sub})$, denoted as PS($Z$) and APS($Z$).
-For estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$, we have PS($Z$) = PS and APS($Z$) = APS because using all six joint levels of ${Z}$ results in the same two strata as above (due to $\pi_{j}(Z)=0$ for some $j=3,4$).
-  3. (iii)
-    The naive estimator. For comparison we include the naive estimator of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, as described in Section 4.1.
-  4. (iv)
-    Estimators within each sub-study.
-For comparison, we also evaluate traditional ANOVA (sample means) and ANCOVA estimators (using linear working models to adjust for $X_{c},X_{b}$, and $Z_{\rm sub}$) based solely on data from each individual sub-study.
-Although these sub-study-specific estimators estimate the treatment effect within each sub-study, which differs from our target estimand defined in (1), we can still compare their variances.
-The simulation results based on $5,000$ runs are reported in Table 2 for the previously described eleven estimators in (i)-(iv).
-The results contain the average bias, standard deviation (SD), the average of the standard error (SE), and the coverage probability (CP) of the $95\%$ confidence interval.
-The true values of estimands are approximated using an independent simulated dataset with size $10^{7}$.
-The following is a summary of the simulation results in Table 2.
-Table 2: Simulation results based on $5,000$ simulation runs
-$\theta_{21}-\theta_{12}=3$ $\theta_{31}-\theta_{13}=1.145$ $\theta_{41}-\theta_{14}=-0.886$
-$n$ Method Bias SD SE CP Bias SD SE CP Bias SD SE CP
-500 Naive -0.231 0.320 0.316 0.874 -0.185 0.342 0.340 0.916 -0.205 0.384 0.380 0.911
-IPW -0.006 0.639 0.636 0.946 0.004 0.776 0.777 0.948 -0.007 0.500 0.497 0.948
-SIPW -0.003 0.341 0.336 0.941 0.005 0.347 0.341 0.943 0.001 0.389 0.381 0.942
-SAIPW -0.018 0.329 0.340 0.951 0.001 0.284 0.284 0.944 -0.001 0.297 0.300 0.949
-$(S)$ 0.000 0.336 0.329 0.939 0.009 0.327 0.324 0.939 0.002 0.356 0.348 0.942
-PS 0.000 0.336 0.335 0.945 0.009 0.327 0.330 0.949 0.002 0.356 0.356 0.946
-APS -0.013 0.329 0.339 0.952 -0.001 0.286 0.289 0.947 -0.002 0.298 0.306 0.956
-PS$(Z)$ Estimate or its SE cannot
-APS$(Z)$ be computed in 408 runs
-ANOVA* 0.001 0.354 0.350 0.946 0.006 0.421 0.417 0.945 0.004 0.424 0.425 0.946
-ANCOVA* -0.008 0.329 0.321 0.942 -0.005 0.319 0.311 0.942 0.002 0.321 0.314 0.942
-1000 Naive -0.230 0.226 0.224 0.819 -0.189 0.240 0.239 0.872 -0.206 0.269 0.268 0.876
-IPW -0.001 0.453 0.451 0.947 0.012 0.550 0.550 0.951 0.003 0.355 0.352 0.943
-SIPW 0.000 0.243 0.239 0.945 0.004 0.246 0.243 0.944 0.001 0.272 0.270 0.948
-SAIPW -0.009 0.232 0.242 0.954 0.004 0.198 0.203 0.955 0.000 0.212 0.213 0.947
-$(S)$ 0.001 0.238 0.234 0.945 0.004 0.233 0.230 0.942 0.003 0.252 0.247 0.943
-PS 0.001 0.238 0.236 0.948 0.004 0.233 0.232 0.944 0.003 0.252 0.250 0.947
-APS -0.006 0.232 0.239 0.952 0.003 0.198 0.203 0.955 0.000 0.213 0.215 0.952
-PS$(Z)$ 0.001 0.238 0.236 0.948
-APS$(Z)$0.005 0.233 0.229 0.942
-ANOVA* 0.003 0.251 0.248 0.947 0.007 0.295 0.294 0.944 0.002 0.301 0.299 0.948
-ANCOVA* -0.003 0.233 0.228 0.947 0.005 0.222 0.220 0.947 0.001 0.227 0.222 0.945
-SAIPW($S$): SAIPW with $\bm{X}=S$.
-PS($Z$) and APS($Z$): post-stratification using all joint levels of $Z=(Z_{\rm win},Z_{\rm sub})$ as strata (PS = PS($Z$) and APS = APS($Z$) for estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$).
-*ANOVA and ANCOVA based on data within each sub-study, estimating the within sub-study treatment effects, which are $3.054$, $1.279$, and $-0.881$ for sub-studies 1-3, respectively.
-  1. 1.
-    The simulation results support our asymptotic theory in Section 5.
-Specifically, the naive estimator is biased, while all IPW-type and PS-type estimators have negligible biases; despite the misspecification of working models, covariate adjustment improves efficiency as reflected by reductions in SD; the PS and SAIPW($S$) estimators demonstrate nearly identical performance.
-Additionally, SEs are close to SDs and CPs are close to the nominal level 95%.
-  2. 2.
-    No results are reported in Table 1 for PS($Z$) and APS($Z$) at $n=500$ because either the estimate of $\theta_{jk}$ or its SE cannot be computed in 408 out of 5,000 simulation runs.
-This issue arises due to insufficient data in the stratum defined by $(Z_{\rm win},Z_{\rm sub})=(3,0)$, where the average number of patients is 10.
-A similar issue occurs in 3 out of 5,000 runs at $n=1,000$, so the results in Table 1 are based on 4,997 runs.
-This problem does not affect the PS and APS estimators: our proposed stratification method combines three strata $(Z_{\rm win},Z_{\rm sub})=(1,0),(2,0),(3,0)$ into ${\cal I}_{12}^{(1)}$ because they all have constant $\pi_{1}(Z)=0.5$ and $\pi_{2}(Z)=0.5$.
-  3. 3.
-    Between IPW and SIPW, it is clear that SIPW is much better in terms of SD under our simulation setting.
-Notably, IPW can have an even larger SD than ANOVA within substudies, which usually does not occur with SIPW.
-  4. 4.
-    The traditional ANOVA and ANCOVA estimators analyze the platform trials as if they were separate, stand-alone studies.
-Without covariate adjustment, the ANOVA estimator has a much larger SD than the SIPW estimator when estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$.
-A similar conclusion applies to the ANCOVA estimator when compared with the SAIPW or APS estimator.
-The efficiency difference is not as pronounced for estimating $\theta_{21}-\theta_{12}$, although the ANOVA and ANCOVA estimators still have larger SDs than the corresponding stabilized IPW-type and PS-type estimators.
-The reason for this is because the traditional ANOVA and ANCOVA do not utilize data across sub-studies, for example, when $n=500$, about 123, 51, and 76 patients are used under treatment arm 1 in sub-studies 1-3, respectively, whereas about 250, 121, and 150 patients in treatment arm 1 under platform trial are shared for our proposed estimators of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, respectively.
-7.2 Real data analysis
-As reviewed in Section 2.2 and shown in Figure 2, the SIMPLIFY design (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study.
-Let $Z$ be the indicator of whether a participant receives HS or DA therapy at baseline and the enrollment window, with the assignment probabilities given in Table 3.
-The enrollment window (EW) is treated as a categorical variable, with EW1 representing the enrollment period before the probability change, and EW2 standing for the enrollment period after the change, as presented in Figure 2.
+Figure 5: Efficiency comparisons among all robust estimators, with the corresponding corollary annotated for each comparison.
+Lastly, we discuss the data used by the estimation methods as an informal way to conceptualize the efficiency comparison.
+When comparing treatments $j$ and $k$, all methods can robustly utilize the outcome data from arms $j$ and $k$ within ${{{\cal I}}}_{jk}$, which may achieve substantial efficiency gains compared to traditional stand-alone trials.
+Furthermore, covariate adjustment methods (including PS, APS, AIPW, and SAIPW) can leverage additional strata and covariate information from all individuals within ${{{\cal I}}}_{jk}$, even those assigned to arms other than $j$ and $k$.
+This can lead to further efficiency gains, especially in platform trials with many arms.
+6 Results
+As reviewed in Section 2.1 and shown in Figure 1, the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study.
+After completing the first study, participants on both HS and DA therapies may re-enroll in the other study contingent on eligibility and willingness to consent.
 The outcome $Y$ is the 6-week mean absolute change in ppFEV1 as described in Section 2.2.
-In our analysis, the treatment effects of interest are $\theta_{21}-\theta_{12}$ and $\theta_{31}-\theta_{13}$.
-Specifically, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS among individuals taking HS, regardless of their DA status.
-This represents the treatment effect among participants with $\pi_{1}(Z)>0$ and $\pi_{2}(Z)>0$, pooled across studies (i.e., participants with $\text{HS}=1$ in Table 3).
-The interpretation applies similarly to $\theta_{31}-\theta_{13}$.
-Note that our estimand is slightly different from the estimand in Mayer-Hamblett et al., (2023), as the latter analyzed the HS and DA studies separately and included participants who re-enrolled for an additional study as noted in Section 2.2.
+In our analysis, the ECE population for evaluating the effect of discontinuing HS versus continuing is the population of all individuals taking only HS or both HS and DA.
+Accordingly, the treatment effect of interest, $\theta_{21}-\theta_{12}$, is defined on this ECE population.
+A similar interpretation applies to $\theta_{31}-\theta_{13}$.
+Demographic and baseline characteristics for participants in the two ECE populations are presented in Tables S1-S2 of the Supplement.
+Note that our estimand differ slightly from those in Mayer-Hamblett et al., (2023).
+They analyzed the HS and DA studies separately according to their primary hypotheses of interest and included data collected after participants re-enrolled in an additional study.
+In contrast, we excluded data recorded after re-enrollment, as our methods currently do not account for repeated measures.
+We plan to address this in future research.
 We also excluded 10 (1.7%) participants due to missing outcomes, resulting in a sample size of $n=584$: 293 in treatment 1, 119 in treatment 2, and 172 in treatment 3.
-Table 3: Assignment probabilities in the SIMPLIFY trial
-${Z}$ Study Treatment
-HS DA EW HS DA $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$
-1 1 EW1 0.50 0.50 0.50 0.25 0.25
-1 1 EW2 0.75 0.25 0.50 0.375 0.125
-1 0 any 1 0 0.50 0.50 0
-0 1 any 0 1 0.50 0 0.50
-We consider the IPW, SIPW, SAIPW, PS, and APS estimators, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section 7.1.
-For SAIPW and APS, in addition to $Z$, we adjust several other baseline covariates, including two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, delta F508 heterozygous, and other or unknown).
-Estimates and 95% confidence intervals are presented in Figure 4, where all values are given in %.
-Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 4), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA.
-The conclusion is consistent with those in Mayer-Hamblett et al., (2023).
-Figure 4: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial
+For the defined estimands, we consider the IPW, SIPW, SAIPW, PS, and APS estimators, as described in Section 4, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section S2 of the Supplement.
+Here, $Z$ includes baseline use of HS, DA, and enrollment windows.
+For SAIPW and APS, $\hat{\mu}_{jk}$ is estimated by fitting a linear regression of $Y$ on covariates using data from treatment arm $j$ in $\mathcal{I}_{jk}$.
+The covariates include $Z$ and several other baseline variables: two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, heterozygous, or other/unknown).
+As shown in Corollary 2 and simulations, when adjusting for covariates using linear working models, the AIPW estimator is numerically nearly identical to the SAIPW estimator and thus omitted.
+We also include the naive ANOVA as described in Section 4.2, which is subject to confounding bias.
+Results are shown in Figure 6.
+Figure 6: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial.
+Naive refers to the naive ANOVA method described in Section 4.2.
+ANOVA* and ANCOVA* are estimates based on data from each individual sub-study.
+IPW, SIPW, SAIPW, PS, and APS are the correct estimators for the estimand defined on the ECE population.
+For all five proposed estimators (IPW, SIPW, SAIPW, PS, and APS), all point estimates are similar, as they target the same estimand defined on the ECE population.
+Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 6), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA – consistent with the findings reported in Mayer-Hamblett et al., (2023).
+IPW and SIPW produce similar confidence intervals, likely because the true effect is near zero in this non-inferiority study.
+SIPW and SAIPW, as well as PS and APS, also yield similar confidence intervals, as most covariates are only weakly associated with the outcome ($R^{2}$ between 0.023 and 0.1 across three arms).
+This likely also explains why the naive estimator does not exhibit substantial bias in this analysis.
+For comparison, we also include the ANOVA* (sample means) and ANCOVA* estimators (using linear working models to adjust for the same covariates included in SAIPW and APS separately for each arm) based solely on data from each individual sub-study.
+As discussed in Section 3.2, these estimators target analysis populations assigned to the HS or DA study.
+As a result, the interpretations of these estimators differ from the five proposed estimators.
+Specifically, for the HS comparison, the HS study includes 18 HS-only and 222 HS+DA participants, while the sample from the ECE population includes 18 HS-only and 427 HS+DA participants.
+For the DA comparison, the DA study includes 139 DA-only and 205 HS+DA participants, whereas the sample from the ECE population includes 139 DA-only and 427 HS+DA participants.
+More detailed demographic and baseline characteristics for participants in the HS and DA studies are presented in Tables S3 and S4.
+The larger population difference in the DA comparison likely explains the more noticeable discrepancy in point estimates for that result (opposite sign relative to the five proposed estimators).
+Moreover, as expected, the ANOVA* and ANCOVA* estimators yield wider confidence intervals compared to the five proposed estimators.
+This is because, for the five proposed estimators, participants on both HS and DA assigned to treatment 1 ($n=214$) contribute to both the treatment 2 vs.
+1 and treatment 3 vs.
+1 comparisons.
+The covariate-adjusted methods (SAIPW and APS) further leverage covariate information from all participants on both HS and DA ($n=427$) in both comparisons.
+In contrast, ANOVA* and ANCOVA* split these participants between the HS and DA sub-studies, reducing the sample size for each comparison.
+As a result, the five proposed estimators achieve approximately 5%–27% variance reduction relative to the sub-study-specific estimators.
+Overall, the results from the SIMPLIFY trial highlight the critical importance of clearly defining the estimand in umbrella and platform trials.
+Without this clarity, estimators may have interpretations that depend on the randomization ratio or trial design.
+An additional advantage of the five proposed estimators is the potential for efficiency gains through robust use of pooled data across sub-studies.
 7 Summary and Recommendations
-Platform trials offer the potential to enhance statistical efficiency by facilitating the sharing of data between different treatment evaluations.
-However, realizing this potential requires careful consideration of several challenges, such as defining meaningful estimands that reflect eligibility and concurrency, and developing robust estimation techniques that accommodate the varying probabilities of treatment assignments over time and differing patient characteristics.
-In this article, we explicitly present how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients.
+Master protocol trials have the potential to enhance statistical efficiency by facilitating the sharing of data across treatment evaluations.
+However, realizing this potential requires careful consideration of two fundamental challenges: defining meaningful estimands to reflect eligibility and concurrency, and developing robust estimation methods to accommodate the treatment assignment probabilities varying with time and individual characteristics.
+In this article, we present a clear framework for constructing a clinically meaningful estimand with precise specification of the population.
+Our main discussion centers on comparing two treatments within the ECE population, as it preserves the integrity of randomization and avoids arbitrary dependence on randomization probabilities.
+This framework is adaptable to various contexts, such as comparisons across multiple treatments and evaluation of treatment effects within specific subgroups or a re-weighted population.
+Achieving clarity on the estimand is crucial for addressing other key issues in future statistical research, including interim analyses and the use of nonconcurrent controls.
 To estimate and infer this estimand with minimal statistical assumptions, we develop methods using weighting or post-stratification to account for varying treatment assignment probabilities.
 We also consider model-assisted approaches for covariate adjustment.
+Theoretical and empirical results are given and a detailed relative efficiency comparison is in Figure 5.
 We provide explicit formulas for robust variance estimation, which have been implemented in our R package, RobinCID.
-Through theoretical analysis and simulation studies assessing the relative efficiency of these methods, we find that the stabilized IPW (SIPW) and post-stratification (PS) significantly enhance the efficiency of the IPW estimator.
-Additionally, model-assisted approaches for covariate adjustment (SAIPW or APS) generally improve efficiency while maintaining robustness against model misspecification.
-A detailed relative efficiency comparison is in Figure 3.
-Note that although SIPW is less efficient than PS asymptotically, it is still useful when there are sparse strata because SIPW may have better finite-sample performance and there may be situations where the PS estimator or its variance estimator cannot be computed.
+Lastly, our results are applicable not only to master protocol trials but also more broadly to trials with constrained or non-uniform treatment assignment.